### v1

# Towards Stable and Faithful Inpainting

Abstract
Recent $*{}^{*}$: Equal contribution. ${}^{\dagger}$: Corresponding author. in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing ill-conditioned problems. However, this enhanced generation often introduces instability, leading to arbitrary object generation within masked regions. This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generative capacity. Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a reconstruction-based masked auto-encoder (MAE) as a stable prior. Aligned with the robust Stable Diffusion inpainting model (SD), ASUKA significantly improves inpainting stability. ASUKA further aligns masked and unmasked regions through an inpainting-specialized decoder, ensuring more faithful inpainting. To validate effectiveness across domains and masking scenarios, we evaluate on MISATO, a collection of several existing dataset. Results confirm ASUKA’s efficacy in both stability and fidelity compared to SD and other inpainting algorithms.
Figure 1: Comparison between ASUKA and reconstruction-based (MAE [20]) as well as generative-based (MAE-FAR [8], MAT [26], Co-Mod [61], LaMa [45], Stable Diffusion Inpainting Model [39]) inpainting algorithms. MAE offers stable masked region estimates, yet falls short in texture detail. GAN-based inpainting struggles with low fidelity, such as omitted white horizontal lines. SD is powerful but unstable, often introducing random elements and suffers from mask-unmask color inconsistency. ASUKA ensures consistency of masked-unmasked areas during diffusion and decoding processes, achieving stable and faithful inpainting. This image is from COCO [30].
1 Introduction
Image inpainting [5] generates content to fill in masked areas of images while maintaining consistency with the unmasked areas. Traditional inpainting algorithms [5, 14, 19, 25, 40] primarily rely on information from unmasked regions to reconstruct masked regions, resulting stable estimations but lacks fidelity. Moreover, the conventional reconstruction loss is prone to generating blurred synthesis due to the inherent ill-posed inpainting issues [36].
To facilitate inpainting with complex mask structures, recent approaches [35, 28, 7, 45, 46, 61] widely utilize the generative models [18, 22, 16], which introduce randomness into the inpainting process, like learning connections between image and noise distributions [18, 22] and utilizing Gibbs sampling [55, 11]. These induced stochasticity alleviates the ill-posed issues, yielding more flexible and faithful generation results [61, 62, 26].
However, every benefit comes with a cost. Current approaches mainly use random masking strategy to train inpainting algorithms, which usually leads to masking foreground objects. When trained with these masks, the model learns to generate foreground objects without information of the object, but sorely based on unrelated background. This introduces a bias in generative models, causing them to hallucinate unreasonable objects in the masked areas. This contradicts the goal of achieving stable inpainting.
In this paper, we propose to combine the strengths of both regression-based reconstruction and distribution-aligned generation to achieve stable and faithful inpainting simultaneously. Particularly, we adopt the Masked Auto-Encoder (MAE) [20] and Stable Diffusion inpainting model (SD) [39], respectively. MAE is trained simply to estimate masked pixels through the L2 reconstruction loss, resulting in stable and average results, albeit with some blurring. On the other hand, SD has impressive generation capacity, but may generate implausible content. In this paper, we demonstrate that the exceptionally faithful feature priors from MAEs can substantially enhance the inpainting stability of SD without sacrificing the performance as shown in Figure 1 with a frozen SD.
Particularly, we manipulate the generation and decoding processes of SD to achieve stable and faithful inpainting with the information from unmasked region.
i) In the generation process, we use MAE to estimate masked region sorely from unmasked region. Specifically, we remove the text condition and turn SD into a MAE prior guided image inpainting model. To align SD with MAE while keeping SD frozen, we propose an alignment module, which is trained on MAE-prior-to-image task with standard diffusion objective and achieves dimension and distribution alignment to cross-attention layers of SD.
ii) In the decoding process, we ensure consistency between the masked and unmasked regions. The use of Variational Auto-Encoder trained with KL divergence loss (KL-VAE) in SD inevitably leads to degradation, causing the color inconsistencies in some inpainting cases. Zhu et al. [64] observed similar issue and train a decoder conditioned on the unmasked regions. However, their decoder only preserves the unmasked region and still suffers from the inconsistency between the masked and unmasked regions. To tackle this challenge, we re-formulate the decoder of KL-VAE as a local harmonization model conditioned on unmasked image pixels. Such a specialized inpainting decoder significantly improves the perceptual consistency of inpainted images.
Combined together, our proposed system, dubbed Aligned Stable inpainting with Unknown Areas prior (ASUKA), is illustrated in Figure 2. To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, we utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30]. This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation. Through experiments conducted on MISATO, especially with large irregular masks, we validate the efficacy of ASUKA.
Contributions. Our contributions can be summarized as:
  • We emphasize stability in inpainting tasks as the generative models suffers from the instability problem;
  • We balance fidelity and stability through aligning reconstruction-based Masked Auto-Encoder (MAE) and generation-based Stable Diffusion (SD).
  • We ensure mask-unmask consistency in the decoding process of SD by re-formulating decoding as a local harmonization task conditioned on unmasked image pixels and training a inpainting-specialized decoder.
2 Related Works
Image inpainting is a longstanding challenge to infill the missing regions with consistent pixels. Classical approaches, relying on patchmatch [15, 4, 58] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes. Benefits from the adversarial training [18], learning-based GANs achieve prominent success in image inpainting [36, 54, 61, 26, 8]. In particular, GAN-based inpainting methods have evolved with adaptive convolutions [31, 54, 57], attention mechanism [53, 52, 56, 24], and high-resolution generalization in frequency learning [45, 51, 13]. Methods like Co-Mod [61] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [26, 62], yielding robust and high-quality generations, even for large masks. However, they suffer from unstable results and uncontrollable hallucinations conditioned on different random latent codes. On the other hand, methods emphasizing higher reconstruction penalties [35, 45, 8] achieve more stable generations, but struggle with large masks, often resulting in blur artifacts. Recent diffusion models [41, 39] have rapidly developed and demonstrated impressive inpainting results. However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction. As a consequence, these diffusion-based approaches face challenges in maintaining faithful inpainting results.
Adapting Stable Diffusion. Currently, SD [39] stands out as one of the most widely used diffusion models. SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space. Various tuning and aligning methods have been developed to introduce additional conditions to SD. This includes image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [17], LoRA fine-tuning [23], and controlnet [59] to add other types of conditions. In this paper, our objective is to enhance stability while preserving the generation capacity of the stable diffusion model. Consequently, we refrain from fine-tuning the U-Net backbone of SD. Conversely, the text-condition is not utilized in inpainting tasks. Textual-inversion [17] introduces learnable tokens in Clip text encoders to improve subject description. ObjectStitch [43] employs the CLIP visual feature to represent the subject. In our approach, we advocate removing the text-condition and instead guiding SD using the Masked Auto-Encoder (MAE) prior for masked region.
Information loss in SD latent space. Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images. OpenAI [2] proposes a larger decoder to improve the decoding quality of SD’s latent. Luo et al. [33] propose a frequency-augmented decoder to address the super-resolution case. Zhu et al. [64] propose to preserve unmasked regions during decoding. In this paper, inspired by these approaches, we take a step further to ensure the low-frequency semantic consistency between mask and unmask regions.
Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning. Typical MIM methods [3, 20, 50, 12] split images into visible and masked patches, learning to estimate masked patches from visible patches. Training targets for visible patches encompass pixel values [20], HOG features [48], and high-level semantic features [49]. While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy. Cao et al. [8] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies. In contrast, this paper proposes facilitating MAE as the inpainting prior for SD and enhancing the stability of the inpainting process by incorporating MAE prior guidance into the cross-attention modules of SD.
3 Methodology
Problem setup. Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region. The target of inpainting is to fill the missing region based on the information of unmasked regions to generate high-fidelity images. In this paper, we focus on the standard inpainting task without utilizing other conditions.
Figure 2: ASUKA adopts the MAE to provide a low-resolution stable prior for frozen SD to maintain the generation capacity while increasing stability. A better decoder is utilized to achieve mask-unmask semantic consistency when decoding SD results from latent to image for better fidelity.
Overview. The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Figure 2. ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39]. Our target is to provide more stable and faithful inpainting results while fully exploiting the generation capacity of SD. Our alignment includes i) a stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a faithful alignment to align ground-truth unmasked region with generated masked region during decoding. To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase stability. To align the MAE prior to SD, we introduce a alignment module, which is trained via the training objective of SD. Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask inconsistency to SD, we train an inpainting-specialized decoder to better decode the latent back to the image space for seamless fidelity. Combined together, ASUKA achieves more stable and faithful inpainting results.
3.1 Masked Auto-Encoder Prior
Figure 3: Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
MAE as stable prior. While recent generative models rely on stochastic noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly. On the other hand, MAE is known to provide a stable estimation of masked regions based purely on the unmasked regions. As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a stable prior for generative models to not generate new concepts. However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the stable diffusion model to perform image-to-image style inpainting, as illustrated in Figure 3. To this end, we adopt the MAE to provide prior to stabilizing SD.
Masking strategy for MAE. However, the original MAE is trained to estimate random masks uniformly distributed in the image. It is infeasible to directly adopt pre-trained MAE for inpainting task, which usually contains large continuous holes in the masked region. Inspired by Cao et al. [8], we fine-tune the MAE to inpainting masks. To adapt MAE for more practical inpainting scenarios, unlike Cao et al. [8] which only use continuous masks of range $[0.1,0.5]$, we construct a systematic masking strategy. Specifically, we generate masks with 50% object-shape mask, 20% Co-Mod mask [61], 20% LaMa [45] mask, 5% rectangle mask and 5% complement rectangle mask. We combine object-shape mask with Co-Mod style and LaMa style with the chance of 25%, respectively. This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks. We control the mask ratio in the range of $[0.1,0.75]$ to follow the training scenario of MAE. For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions.
In practice, we train the MAE on images of size $256^{2}$, which is efficient and produce stable guidance for stable diffusion to generate high-resolution images.
3.2 Align MAE Prior with Stable Diffusion
Replace text-condition with MAE prior. SD is not trained on MAE priors. Fortunately, the text-condition part of SD is not utilized in inpainting, as we do not assume a text condition for inpainting task. Hence, we propose to replace the text-condition part of SD with our proposed MAE prior condition. However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior. To achieve a proper guidance for SD, we introduces the alignment module to align both dimension and distribution.
Dimension alignment. Particularly, we adopt the MAE prior $F_{\mathrm{MAE}}$ of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the MAE feature dimension. To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
Distribution alignment. After aligning the dimension, we use self-attention blocks to learn the distribution alignment for better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD. We train our alignment module with the same masking strategy used to train the MAE, while keeping the continuous-mask fine-tuned MAE and SD frozen. The training objective is the diffusion objective. To encourage the alignment module extract features from MAE prior, in the training stage, we use the full image to generate MAE prior instead of masked image with a chance of $p$.
Figure 4: The illustration of the information loss of KL-VAE. (a) The hue of the image is shifted. (b) KL-VAE still suffers from non-ignorable shifts in low-frequency fields.
3.3 Align Mask-Unmask Regions during Decoding
Information loss of KL-VAE. SD performs all its generative processes in the latent space and subsequently decodes these latent codes back to image space using KL-VAE. Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss. Particularly in tasks like inpainting, the unmasked region necessitates replacement with ground-truth values. Though Rombach et al. [39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Figure 4 (b). Moreover, as shown in Figure 4 (a), repeated decoding with KL-VAE not only noticeably degrades high-frequency details but also shifts in semantic hue. As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant perceptual inconsistencies. This results in low-fidelity inpainting results, especially under the irregular mask structures.
Figure 5: Training pipeline of the decoder, enhancing mask-unmask consistency via augmentations in color and latent spaces.
Figure 6: Qualitative SD1.5 inpainting results decoded by (b) vanilla KL-VAE decoder provided from [39], (c) larger decoder from [64], and (d) our augmented decoder. Our decoder can largely alleviate the mask-unmask color inconsistency.
Unmask-region conditioned decoder. The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes. Inspired by Zhu et al. [64], we design a larger decoder to preserve the global consistency in low-frequency fields. This decoder, as shown in Figure 2, involves incorporating masked images in the pixel-wise color space as additional inputs. Unlike [64], we concatenate the 0-1 mask as the extra input to more precisely locate masked regions. However, the larger decoder with even pixel-wise image inputs still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Figure 6 (c). The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue.
Mask-unmask consistent decoder. To train the decoder to ensure the consistency between generated latent and unmasked pixels, we estimate and enlarge the inconsistency during the training of the decoder. Specifically, we propose the color and latent augmentation as shown in Figure 5 to enhance the decoder to explicitly harmonize masked regions based on unmasked pixels.
i) Color augmentation: all training images would be augmented in brightness, contrast, saturation, and hue. This color augmentation is to avoid the decoder simply learns to lighter the image, as most VAE degradation results in a darker hue and oversaturation.
ii) Latent augmentation: To simulate the real VAE degradation, we propose leveraging the artifacts generated from the diffusion model. However, denoising to real images iteratively is notably time-consuming, even with DDIM [42]. Instead, we propose the one-step estimation of degraded latent $\tilde{\bm{z}}_{0}$ via the encoded clean latent $\bm{z}_{0}$ and noised latent $\bm{z}_{t}$:
$\tilde{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\hat{\bm{z}}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the cumulative production of prescribed variance schedule in diffusion; $\varepsilon_{\theta}(\cdot)$ is the noise estimation from frozen SD based on noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$. Despite $\tilde{\bm{z}}_{0}$ suffering from an extremely large step denoising, we empirically find that the inpainting fine-tuned SD could produce reliable results with little artifacts enhanced by the concatenation of unmasked latent $\bm{z}_{0}$. Moreover, the latent augmentation makes the image undergo two-round encoding and decoding from KL-VAE. This not only facilitates the incorporation of natural VAE degradation into the decoder training but also enhances the overall robustness of the model. We empirically apply 50% to operate the latent augmentation to training images. We fine-tune the decoder based on [64], showcasing superior consistency as compared in Figure 6.
4 Evaluation Dataset
Figure 7: Examples of masked images constructed in MISATO. The first three cases focus on different image domains, while in the last case we focus on the inpainting of background.
To validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30]. Matterport3D contains a indoor image dataset, Flickr-Landscape is consisted of outdoor landscape images, MegaDepth involves a dataset of buildings, and COCO has precise segmentation of foreground subjects. For each dataset, we select 500 representative examples, yielding 2000 testing examples of size $512^{2}$.
The principle of constructing MISATO is to select the most representative and diverse examples. To this end, for Matterport3D, Flickr-Landscape and MegaDepth, we use CLIP visual model [37] to extract semantic visual features. Then we use BisectingKMeans [44] to cluster each dataset into 500 clusters, and select the cluster centers as the evaluation data. The selected data are center cropped and then resized to $512^{2}$. For COCO, we focus on the background inpainting. To this end, for each data we identify the foreground with provided segmentation and remove it from the generated masks, yielding a dataset specified for purely background inpainting.
Combined together, MISATO contains four inpainting domains, indoor, outdoor landscape, building, and background. As shown in Figure 7, for each evaluation image, we adopt the similar masking strategy as in Section 3.1 while not using the rectangle and complement rectangle masks. The masking ratio is set as $[0.2,0.8]$.
5 Experiments
Implementation details. ASUKA utilizes the stable diffusion text-guided inpainting model v1.5 as the generative backbone. We train ASUKA on Places2 [63]. For the MAE [20] used in ASUKA, we fine-tune from the MAE used in Cao et al. [8] with a batch size of 1024. We use four self-attention blocks for the alignment module, training from scratch with AdamW [32] optimizer of learning rate 5e-2 for 23K steps of batch size 64 with the standard diffusion objective. We use warm-up strategy to set initial $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2k training steps and then freeze $p=10\%$. We fine-tune the decoder based on [64] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay. We use ColorJitter to implement color augmentation, with the setting of brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
Evaluation metrics. We compare inpainting algorithms using the following metrics: Peak Signal-to-Noise Ratio (PSNR) to compare the ratio of real data to introduced noise, Structural Similarity Index Measure (SSIM) [47] to identify the perceived change in structural information, Learned Perceptual Image Patch Similarity (LPIPS) [60] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [21] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [61] to measure the human-inspired linear separability.
5.1 Comparison on MISATO
Figure 8: Qualitative comparison on MISATO. We compare ASUKA with Co-Mod [61], MAT [26], LaMa [45], MAE-FAR [8], and SD [39] in landscape (first two rows), indoor (3rd and 4th rows), building (5th and 6th rows) and background inpaintings (last two rows). SD performs unconditional generation. SD-text utilizes text prompt of “background”. SD-prompt utilizes trained prompt for inpainting task using the same training setting of ASUKA. Inpainting algorithms usually suffer from fidelity issue, while SD-based algorithms suffer from stability issue and generate random elements. In contrast, ASUKA achieves stable and faithful inpainting.
[…]
Co-Mod [61] 21.805 0.787 0.179 17.421 0.243 0.109
MAT [26] 21.553 0.787 0.176 17.261 0.255 0.122
LaMa [45] 23.677 0.820 0.155 15.436 0.260 0.135
MAE-FAR [8] 24.150 0.821 0.142 13.283 0.282 0.153
SD [39] 22.454 0.796 0.168 12.812 0.345 0.211
SD-text 22.394 0.793 0.164 12.603 0.337 0.207
SD-prompt 22.381 0.798 0.160 12.517 0.331 0.204
ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
Table 1: Quantitative comparison on MISATO. Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8] are state-of-the-art inpainting algorithms. SD [39] performs unconditional generation. SD-text utilizes text prompts of “background” to perform text-guided generation. SD-prompt utilizes learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation. ASUKA and SD use the stable diffusion text-guided inpainting model v1.5.
Competitors. We first compare ASUKA with SD [39]. Specifically, we consider three variants of SD: i) SD, which generates inpainting results with no guidance, and performs unconditional generation; ii) SD-text, which generates inpainting results guided by the text prompt of “background”; iii) SD-prompt, where we replace the text prompt with learnable tokens and train the tokens using the same training pipeline of ASUKA. Beyond SD, we also compare ASUKA with state-of-the-art inpainting algorithms, including Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8].
Quantitative comparison results are reported in Table 1. ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS. Particularly, the U-IDS and P-IDS coincide with human preference [61], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin. The inconsistency results between PSNR, SSIM, LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net. On the other hand, the PSNR and SSIM is known to have bad correlation with human preferences, as several illustrative examples shown in the appendix. In summary, the results validate the contribution of ASUKA that yields stable and faithful inpainting through frozen SD.
Qualitative comparison examples are shown in Figure 8. i) The state-of-the-art inpainting algorithms usually suffer from fidelity issues, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row. LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks. ii) The three variants of SD usually suffer from the stability issue and generate random new elements in the masked region, in almost all the illustrated images. iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows. On the other hand, SD generates random elements in the masked region. iv) In contrast, ASUKA enjoys superior stable and faithful inpainting.
5.2 Further Analysis of ASUKA
[…]
p.t. MAE 22.485 0.799 0.151 11.513 0.354 0.258
f.t. MAE 22.495 0.799 0.150 11.460 0.368 0.256
Table 2: Quantitative comparison of ASUKA using pre-trained MAE (p.t. MAE) from [8] and fine-tuned MAE (f.t. MAE).
Ablation of MAE prior. We compare our fine-tuned MAE with directly adopting the MAE trained in [8] without further fine-tuning. To this end, we train ASUKA with the MAE in [8] using the same training strategy and compare the results in. Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS. This improvement comes from the better adaptation on the real-world masks.
[…]
linear 22.402 0.797 0.155 11.934 0.361 0.227
attn 22.474 0.798 0.152 11.613 0.362 0.234
cross x4 22.488 0.799 0.152 11.762 0.368 0.238
ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
Table 3: Ablation of different alignment modules. Linear adopts linear layer to align feature dimension only; attn adopts a self-attention layer to further align distribution; cross x4 adopts 4 cross-attention layers; while ASUKA adopts 4 self-attention layers.
Ablation of alignment module. Our alignment module includes a dimension and a distribution alignment. We validate the effectiveness of each module by comparing with the following variants step by step: i) linear: Use linear layer to align feature dimension only; ii) attn: Based on i, further use a single self-attention block to align the distribution; iii) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior. ASUKA adopts 4 self-attention blocks. Results are shown in Table 3. The self-attention block shows improved results on simply aligning the dimension as well as using a cross-attention block to align the distribution. Using 4 self-attention blocks improves the capacity of the alignment module.
[…]
KL-VAE [39] 22.454 0.795 0.156 11.949 0.343 0.208
Asy. VAE [64] 22.566 0.801 0.151 11.634 0.361 0.231
color-aug 22.475 0.799 0.152 11.603 0.357 0.229
Ours 22.495 0.799 0.150 11.460 0.368 0.256
Table 4: Quantitative comparison of different decoders for SD.
Ablation of decoder training pipeline. For the decoder, we compare ASUKA with i) KL-VAE: the decoder used in SD; ii) Asy. VAE: the decoder trained by [64]; iii) color-aug: only use color augmentation to train the decoder; Results are shown in Table 4, which shows the superiority of our augmented decoder to ensure the mask-unmask consistency.
Figure 9: The curse of self-attention, causing the MAE falsely estimate the masked region and powerful text-guided diffusion models fail to generation content based on text prompts. ASUKA potential circumvents this issue by using a blank paper image as the input to the MAE to provide correct prior.
Limitation: The “curse” of self-attention. The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module. Specifically, as showin in Figure 9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal. Notably, this curse of self-attention significantly impacts diffusion-based generative models. It results in the inability to accurately follow “blank paper” text prompts, even when employing a substantial classifier-free guidance scale of 9. This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [38] and Adobe’s FileFly [1]. Nevertheless, ASUKA has the potential to circumvent this issue by modifying the MAE prior, for instance, by instead using a blank paper image as the input to MAE prior. A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
6 Conclusion
In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve stable and faithful inpainting via frozen Stable Diffusion inpainting model (SD). To achieve stability, we adopt a reconstruction-based masked-auto encoder (MAE) as the stable prior for masked region purely from unmasked region. Then we align the stable prior to frozen SD with the proposed alignment module. To achieve fidelity, we resolve the mask-unmask inconsistency of SD in the decoding process. We fine-tune a unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process. To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets. ASUKA enjoys superior stable and faithful inpainting results than state-of-the-art inpainting algorithms and SD-based variants.
Supplementary Material
7 Extra Analysis
7.1 Comparison on Places2
[…]
Co-Mod [61] 19.413 0.709 0.267 5.794 0.274 0.096
MAT [26] 19.729 0.740 0.202 3.765 0.348 0.195
LaMa [45] 21.503 0.771 0.202 6.693 0.247 0.050
MAE-FAR [8] 22.140 0.779 0.174 3.559 0.307 0.105
SD [39] 20.302 0.746 0.193 1.514 0.375 0.207
SD-text 20.359 0.748 0.191 1.506 0.373 0.202
SD-prompt 20.548 0.753 0.189 1.477 0.390 0.234
ASUKA 20.406 0.752 0.183 1.230 0.413 0.287
Table 5: Quantitative comparison on Places 2.
We evaluation ASUKA and other inpainting algorithms on the standard testing benchmark dataset of Places2 [63], which includes 36500 testing examples. Results on this dataset also validate the efficacy of ASUKA.
7.2 User-Study
Model Consistency (%) Quality (%)
Co-Mod [61] 4.69 5
MAT [26] 7.5 3.33
LaMa [45] 8.44 8.33
MAE-FAR [8] 5.31 5.83
SD [39] 10.63 5.83
SD-text 7.81 15.83
SD-prompt 16.25 15.83
ASUKA 39.38 40
Table 6: User-study of top-1 ratio on consistency and quality among all the inpainting results.
To evaluate the user preference on inpainting algorithms, we conduct an user-study. Specifically, we randomly select 40 testing images along with inpainting results on these images. We ask the user to select the best inpainting results from the following perspectives respectively: i) Consistency: the masked region should be consistent with unmasked region, involving color and texture consistency and not generating new elements; ii) Quality: the fidelity of the inpainting results. The first metric measures the stability while the other measures the fidelity of the inpainting results. We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms. This result validate the efficacy of ASUKA on alignment with human preference.
7.3 The Misalignment Between PSNR, SSIM and Human Preference
The PSNR and SSIM are known to be have bad correlation with human preferences. To show this, we report the PSNR and SSIM scores for the qualitative comparison on MISATO ( Figure 8 in the paper) in Figure 10. Higher PSNR and SSIM do not always indicate better results that human prefers. Particularly, blurred inpainting results usually have higher PSNR and SSIM scores, like the 1st, 2rd, 4th, 8th rows. However, it is clear that blurred results are not expected.
Figure 10: Misalignment between PSNR, SSIM and human preference. Images are same with Figure 8. PSNR and SSIM scores are listed in the bottom, where the best values are marked as red.

### v2

@@
-Towards Stable and Faithful Inpainting
+Project Page: https://yikai-wang.github.io/asuka/ MISATO Dataset: https://github.com/Yikai-Wang/asuka-misato
 Abstract
-Recent $*{}^{*}$: Equal contribution.
-${}^{\dagger}$: Corresponding author.
-in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing ill-conditioned problems.
-However, this enhanced generation often introduces instability, leading to arbitrary object generation within masked regions.
-This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generative capacity.
-Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a reconstruction-based masked auto-encoder (MAE) as a stable prior.
-Aligned with the robust Stable Diffusion inpainting model (SD), ASUKA significantly improves inpainting stability.
-ASUKA further aligns masked and unmasked regions through an inpainting-specialized decoder, ensuring more faithful inpainting.
-To validate effectiveness across domains and masking scenarios, we evaluate on MISATO, a collection of several existing dataset.
-Results confirm ASUKA’s efficacy in both stability and fidelity compared to SD and other inpainting algorithms.
-Figure 1: Comparison between ASUKA and reconstruction-based (MAE [20]) as well as generative-based (MAE-FAR [8], MAT [26], Co-Mod [61], LaMa [45], Stable Diffusion Inpainting Model [39]) inpainting algorithms.
-MAE offers stable masked region estimates, yet falls short in texture detail.
-GAN-based inpainting struggles with low fidelity, such as omitted white horizontal lines.
-SD is powerful but unstable, often introducing random elements and suffers from mask-unmask color inconsistency.
-ASUKA ensures consistency of masked-unmasked areas during diffusion and decoding processes, achieving stable and faithful inpainting.
-This image is from COCO [30].
+Recent progress in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing large irregular masks.
+However, this enhanced generation often introduces context-instability, leading to arbitrary object generation within masked regions.
+This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generation capacity.
+Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a Masked Auto-Encoder (MAE) to produce reconstruction-based prior.
+Aligned with the powerful Stable Diffusion inpainting model (SD), ASUKA significantly improves context stability.
+ASUKA further adopts an inpainting-specialized decoder, highly reducing the color inconsistency issue of SD and thus ensuring more visual-consistent inpainting.
+We validate effectiveness of inpainting algorithms on benchmark dataset Places 2 and a collection of several existing datasets, dubbed MISATO, across diverse domains and masking scenarios.
+Results on these benchmark datasets confirm ASUKA’s efficacy in both context-stability and visual-consistency compared to SD and other inpainting algorithms.
+Keywords:
+Image Inpainting Context-Stability Visual-Consistency
+Figure 1: Comparison on $1024^{2}$ image (depicted from MegaDepth [26]) between our proposed ASUKA and reconstruction-based (MAE [19]) as well as generative-based (MAE-FAR [7], MAT [25], Co-Mod [59], LaMa [45], Stable Diffusion [39]) inpainting models.
+MAE offers stable estimates, yet falls short in texture detail.
+GAN-based inpainting struggles with low fidelity.
+Powerful SD hallucinates random elements and suffers from mask-unmask visual inconsistency (color of clouds in the orange square).
+ASUKA achieves context-stable and visual-consistent inpainting.
 1 Introduction
-Image inpainting [5] generates content to fill in masked areas of images while maintaining consistency with the unmasked areas.
-Traditional inpainting algorithms [5, 14, 19, 25, 40] primarily rely on information from unmasked regions to reconstruct masked regions, resulting stable estimations but lacks fidelity.
-Moreover, the conventional reconstruction loss is prone to generating blurred synthesis due to the inherent ill-posed inpainting issues [36].
-To facilitate inpainting with complex mask structures, recent approaches [35, 28, 7, 45, 46, 61] widely utilize the generative models [18, 22, 16], which introduce randomness into the inpainting process, like learning connections between image and noise distributions [18, 22] and utilizing Gibbs sampling [55, 11].
-These induced stochasticity alleviates the ill-posed issues, yielding more flexible and faithful generation results [61, 62, 26].
-However, every benefit comes with a cost.
-Current approaches mainly use random masking strategy to train inpainting algorithms, which usually leads to masking foreground objects.
-When trained with these masks, the model learns to generate foreground objects without information of the object, but sorely based on unrelated background.
-This introduces a bias in generative models, causing them to hallucinate unreasonable objects in the masked areas.
-This contradicts the goal of achieving stable inpainting.
-In this paper, we propose to combine the strengths of both regression-based reconstruction and distribution-aligned generation to achieve stable and faithful inpainting simultaneously.
-Particularly, we adopt the Masked Auto-Encoder (MAE) [20] and Stable Diffusion inpainting model (SD) [39], respectively.
-MAE is trained simply to estimate masked pixels through the L2 reconstruction loss, resulting in stable and average results, albeit with some blurring.
-On the other hand, SD has impressive generation capacity, but may generate implausible content.
-In this paper, we demonstrate that the exceptionally faithful feature priors from MAEs can substantially enhance the inpainting stability of SD without sacrificing the performance as shown in Figure 1 with a frozen SD.
-Particularly, we manipulate the generation and decoding processes of SD to achieve stable and faithful inpainting with the information from unmasked region.
-i) In the generation process, we use MAE to estimate masked region sorely from unmasked region.
-Specifically, we remove the text condition and turn SD into a MAE prior guided image inpainting model.
-To align SD with MAE while keeping SD frozen, we propose an alignment module, which is trained on MAE-prior-to-image task with standard diffusion objective and achieves dimension and distribution alignment to cross-attention layers of SD.
-ii) In the decoding process, we ensure consistency between the masked and unmasked regions.
-The use of Variational Auto-Encoder trained with KL divergence loss (KL-VAE) in SD inevitably leads to degradation, causing the color inconsistencies in some inpainting cases.
-Zhu et al.
-[64] observed similar issue and train a decoder conditioned on the unmasked regions.
-However, their decoder only preserves the unmasked region and still suffers from the inconsistency between the masked and unmasked regions.
-To tackle this challenge, we re-formulate the decoder of KL-VAE as a local harmonization model conditioned on unmasked image pixels.
-Such a specialized inpainting decoder significantly improves the perceptual consistency of inpainted images.
-Combined together, our proposed system, dubbed Aligned Stable inpainting with Unknown Areas prior (ASUKA), is illustrated in Figure 2.
-To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, we utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30].
+Figure 2: Object removal on $1024^{2}$ images. ASUKA enjoys context-stable (first to second rows) and visual-consistent (second to third rows) inpainting compared with SD.
+Image inpainting [5] fills masked areas of images while maintaining consistency with the unmasked regions.
+Traditional inpainting algorithms [5, 13, 18, 24, 40] often result in blurred synthesis when reconstructing masked regions, as indicated by [36].
+The advent of Generative Adversarial Networks (GANs) has led to works [35, 27, 8, 45, 46, 59, 17, 21, 15] that utilize GANs to fill complex mask structures, achieving impressive inpainting results.
+However, they still struggle with general challenging inpainting cases, particularly in filling large holes.
+The recent Stable Diffusion Inpainting (SD) [39] model, with its large model capacity and extensive dataset training, acts as a versatile tool for image inpainting.
+However, we have observed a phenomenon of "context-instability" problem, where the model generates random, unreasonable elements to fill masked regions of images, as depicted in first to second rows in Fig.
+2, as well as Fig.
+1 and Fig.
+8.
+This issue appears to come from the random masking strategy used to train SD.
+This strategy introduces training cases where foreground objects are completely masked and the SD is forced to fill masked regions with foreground objects.
+Consequently, SD will hallucinate unreasonable objects devoid of contextual information.
+Adjusting prompts may reduce this risk, but the best prompt is image-dependent, making it infeasible for practical applications.
+Furthermore, the inpainted results of SD may suffer from "visual inconsistency" problem.
+This problem, less explored in academia but critical for real-world industrial applications, results in significant color discrepancies between inpainted and unmasked regions and exhibits smear-like traces in the image, as shown in the second to third rows in Fig.
+2, as well as Fig.
+1 and Fig.
+8.
+Essentially, this visual-inconsistency comes from the misalignment between the pixel distributions of filled results and original images due to information loss in the Kullback-Leibler Variational Autoencoder (KL-VAE) used in SD, as illustrated in Fig.
+5.
+To improve visual consistency, an inpainting-specialized VAE decoder is necessary to explicitly mitigate information loss based on available ground-truth information of unmasked images.
+To efficiently tackle context-instability and visual inconsistency issues, we enhance the SD [39] with regression-based reconstruction and distribution-aligned generation.
+This results in an improved image inpainting model that avoids generating unreasonable elements in the masked region and reduces mask-unmask color inconsistency.
+The vanilla SD [39] adopts a KL-VAE to compress image into latent and perform diffusion in the latent space.
+We manipulate its generation and decoding processes for better context stability and visual consistency.
+i) We propose using the Masked Auto-Encoder (MAE)[19] as a prior to guide and stabilize the SD’s generation process.
+As shown in Fig.
+2, MAE yields stable yet blurred results, while SD may produce implausible content despite its impressive generation capacity.
+By aligning MAE prior with SD objectives, we improve context-stability without compromising performance.
+ii) We redesign the KL-VAE decoder in SD to address color inconsistencies between masked and unmasked regions.
+The new decoder, acting as a local harmonization model conditioned on unmasked image pixels, significantly enhances visual consistency.
+Formally, to enhance the context-stability and visual-consistency of image inpainting model, we present the Aligned Stable Inpainting with Unknown Areas prior (ASUKA) framework.
+i) To enhance context-stability, we replace the text-condition part of the SD with the MAE prior.
+We fine-tune the MAE to ensure adaptability to practical inpainting scenarios, as MAE is trained on uniformly distributed masks but inpainting usually contains large continuous masks.
+We train an alignment module to align MAE prior with SD using the standard SD objective while freezing the fine-tuned MAE and SD.
+ii) To address visual-consistency issues during decoding, we redesign the decoder of the KL-VAE in SD.
+We inject additional inputs of unmasked image pixels and 0-1 masks to the decoder.
+This new decoder is trained to perform local harmonization task, taking as inputs of inharmonious latent and taking as outputs the original harmonious image.
+We generate inharmonious latent via the proposed color and latent augmentation techniques to incorporate KL-VAE information loss and generation artifacts of SD.
+Then the decoder is trained to recover the original image from corrupted latent based on the ground-truth information of unmasked image.
+These steps collectively enable ASUKA to achieve more context-stable and visual-consistent inpainting results.
+To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, in addition to the benchmark dataset Places 2 [61], we further utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29].
 This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation.
-Through experiments conducted on MISATO, especially with large irregular masks, we validate the efficacy of ASUKA.
-Contributions. Our contributions can be summarized as:
-  • We emphasize stability in inpainting tasks as the generative models suffers from the instability problem;
-  • We balance fidelity and stability through aligning reconstruction-based Masked Auto-Encoder (MAE) and generation-based Stable Diffusion (SD).
-  • We ensure mask-unmask consistency in the decoding process of SD by re-formulating decoding as a local harmonization task conditioned on unmasked image pixels and training a inpainting-specialized decoder.
+Experiments on MISATO and Places 2, especially with large irregular masks, validate the efficacy of ASUKA.
+Contributions.
+ASUKA enhances image inpainting from context-stability and visual-consistency while leveraging the generation capacity of the frozen SD inpainting model.
+It achieves this through two main components: i) Context-Stable Alignment: ASUKA aligns the stable MAE prior with SD to provide a context-stable estimation of masked regions, replacing the text-condition part of SD with MAE prior.
+ii) Visual-Consistent Alignment: ASUKA re-formulates the decoding from SD latent to image as a local harmonization task, trains an inpainting-specialized decoder to align masked and unmasked regions during decoding and thus mitigates color inconsistencies caused by information loss from the SD.
 2 Related Works
@@
 Classical approaches, relying on patchmatch [14, 4, 56] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes.
-Benefits from the adversarial training [18], learning-based GANs achieve prominent success in image inpainting [36, 54, 61, 26, 8].
-In particular, GAN-based inpainting methods have evolved with adaptive convolutions [31, 54, 57], attention mechanism [53, 52, 56, 24], and high-resolution generalization in frequency learning [45, 51, 13].
+GAN-based [17] inpainting methods [36, 53, 59, 25, 7] have evolved with adaptive convolutions [30, 53, 54], attention mechanism [52, 51, 55, 23], and high-resolution generalization in frequency learning [45, 50, 12].
 Methods like Co-Mod [59] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [25, 60], yielding robust and high-quality generations, even for large masks.
@@
 However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction.
-As a consequence, these diffusion-based approaches face challenges in maintaining faithful inpainting results.
+As a consequence, diffusion-based approaches also suffer from context-instability issue.
 Adapting Stable Diffusion.
-Currently, SD [39] stands out as one of the most widely used diffusion models.
+SD [39] is one of the most widely used diffusion models.
 SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space.
-Various tuning and aligning methods have been developed to introduce additional conditions to SD.
-This includes image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [17], LoRA fine-tuning [23], and controlnet [59] to add other types of conditions.
-In this paper, our objective is to enhance stability while preserving the generation capacity of the stable diffusion model.
+Various tuning and aligning methods have been developed to introduce additional conditions to SD, including image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [16], LoRA fine-tuning [22], and controlnet [57] to add other types of conditions.
+Our objective is to enhance context-stability while preserving the generation capacity of SD.
 Consequently, we refrain from fine-tuning the U-Net backbone of SD.
-Conversely, the text-condition is not utilized in inpainting tasks.
+The text-condition is not utilized in inpainting tasks.
 Textual-inversion [16] introduces learnable tokens in Clip text encoders to improve subject description.
@@
 Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images.
+See Fig.
+5(a) for illustrative examples.
 OpenAI [2] proposes a larger decoder to improve the decoding quality of SD’s latent.
@@
 [62] propose to preserve unmasked regions during decoding.
-In this paper, inspired by these approaches, we take a step further to ensure the low-frequency semantic consistency between mask and unmask regions.
+In this paper, we ensure the low-frequency color visual-consistency in the decoding process of SD.
 Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning.
@@
 [7] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies.
-In contrast, this paper proposes facilitating MAE as the inpainting prior for SD and enhancing the stability of the inpainting process by incorporating MAE prior guidance into the cross-attention modules of SD.
+In contrast, this paper uses MAE prior to enhance the context-stability of SD.
 3 Methodology
@@
 In this paper, we focus on the standard inpainting task without utilizing other conditions.
-Figure 2: ASUKA adopts the MAE to provide a low-resolution stable prior for frozen SD to maintain the generation capacity while increasing stability.
-A better decoder is utilized to achieve mask-unmask semantic consistency when decoding SD results from latent to image for better fidelity.
+We adopt stable diffusion inpainting model [39] and focus on the general issues of inpainting models, i) context-instability issue: unstable and uncontrollable hallucinations, yielding random elements generated in the masked region; ii) visual-inconsistency issue: mask-unmask color inconsistency issue, yielding smear-like traces in the masked region.
+Figure 3: ASUKA tackles the context-instability issue by adopting the MAE to provide a stable prior for frozen SD to maintain the generation capacity while increasing context-stability.
+For the visual-inconsistency issue, ASUKA utilizes an inpainting-specialized decoder to achieve mask-unmask color consistency when decoding SD latent.
+(a) Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
+(b) Alignment module.
+Figure 4: (a) Use MAE as initial value v.s. as condition. (b) Alignment module.
 Overview.
-The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Figure 2.
+The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Fig.
+3.
 ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39].
-Our target is to provide more stable and faithful inpainting results while fully exploiting the generation capacity of SD.
-Our alignment includes i) a stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a faithful alignment to align ground-truth unmasked region with generated masked region during decoding.
-To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase stability.
-To align the MAE prior to SD, we introduce a alignment module, which is trained via the training objective of SD.
-Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask inconsistency to SD, we train an inpainting-specialized decoder to better decode the latent back to the image space for seamless fidelity.
-Combined together, ASUKA achieves more stable and faithful inpainting results.
+Our target is to provide more context-stable and visual-consistent inpainting results while fully exploiting the generation capacity of SD.
+ASUKA includes i) a context-stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a visual-consistent alignment to align ground-truth unmasked region with generated masked region during decoding.
+To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase context-stability.
+To align the MAE prior to SD, we introduce an alignment module, which is trained via the training objective of SD.
+Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask color inconsistency, we train an inpainting-specialized decoder to decode the latent back to the image space for seamless visual-consistency.
+Combined together, ASUKA achieves more context-stable and visual-consistent inpainting.
 3.1 Masked Auto-Encoder Prior
-Figure 3: Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
-MAE as stable prior.
-While recent generative models rely on stochastic noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly.
-On the other hand, MAE is known to provide a stable estimation of masked regions based purely on the unmasked regions.
-As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a stable prior for generative models to not generate new concepts.
-However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the stable diffusion model to perform image-to-image style inpainting, as illustrated in Figure 3.
+MAE as context-stable prior.
+While recent generative models rely on random noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly.
+In contrast, MAE is known to provide a context-stable estimation of masked regions based purely on the unmasked regions.
+As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a context-stable prior for generative models to not generate new concepts.
+However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the SD to perform inpainting in image-to-image style, as illustrated in Fig.
+3(a).
 To this end, we adopt the MAE to provide prior to stabilizing SD.
-Masking strategy for MAE.
-However, the original MAE is trained to estimate random masks uniformly distributed in the image.
-It is infeasible to directly adopt pre-trained MAE for inpainting task, which usually contains large continuous holes in the masked region.
+Training MAE.
+The original MAE is trained to estimate random masks uniformly distributed in the image, while inpainting task usually contains large continuous masks.
 Inspired by Cao et al.
 [7], we fine-tune the MAE to inpainting masks.
-To adapt MAE for more practical inpainting scenarios, unlike Cao et al.
-[8] which only use continuous masks of range $[0.1,0.5]$, we construct a systematic masking strategy.
-Specifically, we generate masks with 50% object-shape mask, 20% Co-Mod mask [61], 20% LaMa [45] mask, 5% rectangle mask and 5% complement rectangle mask.
-We combine object-shape mask with Co-Mod style and LaMa style with the chance of 25%, respectively.
+To adapt MAE for more practical inpainting scenarios, we construct a systematic masking strategy.
+The mask basis contains: object-shape mask, irregular mask, and regular mask.
+We collect object-shape masks from COCO [29] object segments.
+We use irregular masks from previous studies, including Co-Mod mask [59] and LaMa [45] mask.
+The regular masks contains rectangle and complement rectangle mask.
+To ensure generalization and coverage, for each mask we generate from mask basis with the probability of 50% object-shape, 40% irregular, and 10% regular.
+For object-shape mask basis, we combine it with irregular mask with the chance of 50%.
 This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks.
@@
 For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions.
-In practice, we train the MAE on images of size $256^{2}$, which is efficient and produce stable guidance for stable diffusion to generate high-resolution images.
-3.2 Align MAE Prior with Stable Diffusion
+This benefits ASUKA to tackle the large hole inpainting task.
+3.2 Align MAE Prior with SD for Context-Stability
 Replace text-condition with MAE prior.
@@
 However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior.
-To achieve a proper guidance for SD, we introduces the alignment module to align both dimension and distribution.
+To achieve a proper guidance for SD, we introduce the alignment module to align MAE with SD in both dimension and distribution perspective, as shown in 3(b).
 Dimension alignment.
-Particularly, we adopt the MAE prior $F_{\mathrm{MAE}}$ of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the MAE feature dimension.
+Particularly, the MAE prior $F_{\mathrm{MAE}}$ is of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the feature dimension.
 To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
 Distribution alignment.
-After aligning the dimension, we use self-attention blocks to learn the distribution alignment for better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD.
-We train our alignment module with the same masking strategy used to train the MAE, while keeping the continuous-mask fine-tuned MAE and SD frozen.
-The training objective is the diffusion objective.
-To encourage the alignment module extract features from MAE prior, in the training stage, we use the full image to generate MAE prior instead of masked image with a chance of $p$.
+After aligning the dimension, we use self-attention blocks to learn to better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD.
+We train our alignment module using the standard diffusion objective with the same masking strategy used to train the MAE, keeping other modules frozen.
+Misalignment.
+When training the alignment module with the set (input image, MAE prior, SD result), misalignment issues may arise.
+For example, if an object is completely masked, the MAE will predict the masked area with background, whereas the SD is trained to recreate the object.
+This difference can lead the alignment module to mistakenly disregard the MAE prior.
+To address this, we improve the SD’s adherence to the MAE prior by substituting the MAE predicted prior with the MAE reconstructed prior at a probability of $p$.
+The MAE reconstructed prior involves using MAE to recreate the image without masking any area, ensuring MAE has access to all information needed for reconstruction.
+This approach helps train the alignment module to better guide the SD.
 Figure 5: The illustration of the information loss of KL-VAE. (a) The hue of the image is shifted. (b) KL-VAE still suffers from non-ignorable shifts in low-frequency fields.
-3.3 Align Mask-Unmask Regions during Decoding
+(a) Qualitative SD1.5 inpainting results decoded by (b) vanilla decoder of SD [39], (c) larger decoder [62], (d) our decoder. Our decoder largely alleviate the mask-unmask color inconsistency.
+(b) Decoder trained by local harmonization task, enhancing mask-unmask consistency by reconstructing original image guided by the unmasked region from augments in color and latent spaces.
+Figure 6: ASUKA trains a specialized-decoder to solve the color-inconsistency issue from the information loss of KL-VAE and domain gap between generated and real latent.
+3.3 Mask-Unmask Align during Decoding for Visual-Consistency
 Information loss of KL-VAE.
@@
 Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss.
-Particularly in tasks like inpainting, the unmasked region necessitates replacement with ground-truth values.
+Particularly in tasks like inpainting, we have ground-truth values for the unmasked region.
 Though Rombach et al.
-[39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Figure 4 (b).
-Moreover, as shown in Figure 4 (a), repeated decoding with KL-VAE not only noticeably degrades high-frequency details but also shifts in semantic hue.
-As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant perceptual inconsistencies.
-This results in low-fidelity inpainting results, especially under the irregular mask structures.
-Figure 5: Training pipeline of the decoder, enhancing mask-unmask consistency via augmentations in color and latent spaces.
-Figure 6: Qualitative SD1.5 inpainting results decoded by (b) vanilla KL-VAE decoder provided from [39], (c) larger decoder from [64], and (d) our augmented decoder.
-Our decoder can largely alleviate the mask-unmask color inconsistency.
+[39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Fig.
+5 (b).
+The KL-VAE will not only noticeably degrades high-frequency details but also shifts in semantic hue.
+This shift can be verified by repeated reconstruction with KL-VAE, as shown in Fig.
+5 (a) where larger shift is observed during repeated reconstruction.
+As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant visual-inconsistencies.
+This visual-inconsistency issue is more severe under the irregular or large mask structures.
 Unmask-region conditioned decoder.
 The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes.
-Inspired by Zhu et al.
-[64], we design a larger decoder to preserve the global consistency in low-frequency fields.
-This decoder, as shown in Figure 2, involves incorporating masked images in the pixel-wise color space as additional inputs.
-Unlike [64], we concatenate the 0-1 mask as the extra input to more precisely locate masked regions.
-However, the larger decoder with even pixel-wise image inputs still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Figure 6 (c).
+Zhu et al.
+[62] adopts a larger decoder, with additional inputs of masked images.
+However, it still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Fig.
+5(a) (c).
 The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue.
-Mask-unmask consistent decoder.
-To train the decoder to ensure the consistency between generated latent and unmasked pixels, we estimate and enlarge the inconsistency during the training of the decoder.
-Specifically, we propose the color and latent augmentation as shown in Figure 5 to enhance the decoder to explicitly harmonize masked regions based on unmasked pixels.
-i) Color augmentation: all training images would be augmented in brightness, contrast, saturation, and hue.
-This color augmentation is to avoid the decoder simply learns to lighter the image, as most VAE degradation results in a darker hue and oversaturation.
-ii) Latent augmentation: To simulate the real VAE degradation, we propose leveraging the artifacts generated from the diffusion model.
+To solve the mask-unmask visual-inconsistency issue, we reformulate the decoding as a local harmonization task to preserve the visual consistency.
+As shown in Fig.
+3, compared with SD decoder, our decoder involves additional inputs of masked images in the pixel-wise color space and the 0-1 mask.
+Mask-unmask visual-consistent decoder.
+To train the decoder to ensure visual-consistency between generated latent and unmasked pixels, we re-formulate the decoding as a local harmonization task.
+Specifically, we propose the color and latent augmentation as shown in Fig.
+5(b) to estimate and enlarge the visual-inconsistency during decoding.
+We follow the standard KL-VAE training pipeline of SD, but replacing the inputs with augmented ones.
+Particularly, we use the original image as the reconstruction target and use color and latent augmentation to corrupt input image, simulating the information loss of KL-VAE and domain gap between generated and real latent, respectively.
+This forces the decoder to reconstruct the un-corrupted image based on the ground-truth unmasked region.
+i) Color augmentation: The first information loss comes from the imperfect VAE reconstruction.
+VAE is trained to encode and decode the input image for reconstruction, but suffers from color shift as shown in Fig.
+5(a) (b).
+Empirically, further conditioned on unmasked image alleviate but not solve the color inconsistency issue, as shown in Fig.
+5(a) (c).
+Hence, we need to explicitly train the decoder to ensure color consistency.
+To this end, we augment all training images in brightness, contrast, saturation, and hue, and requires the decoder to reconstruct original image conditioned on the unaugmented unmasked image.
+This encourages the decoder to fully utilize the information from unmasked regions.
+ii) Latent augmentation: The other information loss comes from the distribution gap between SD generated latent and real latent.
+This gap also causes color inconsistency even if we alleviate the VAE reconstruction loss, as shown in Fig.
+6(a).
+To simulate this loss, we incorporate the artifacts generated from the SD to train the decoder.
 However, denoising to real images iteratively is notably time-consuming, even with DDIM [42].
-Instead, we propose the one-step estimation of degraded latent $\tilde{\bm{z}}_{0}$ via the encoded clean latent $\bm{z}_{0}$ and noised latent $\bm{z}_{t}$:
-$\tilde{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\hat{\bm{z}}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
-where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the cumulative production of prescribed variance schedule in diffusion; $\varepsilon_{\theta}(\cdot)$ is the noise estimation from frozen SD based on noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$.
-Despite $\tilde{\bm{z}}_{0}$ suffering from an extremely large step denoising, we empirically find that the inpainting fine-tuned SD could produce reliable results with little artifacts enhanced by the concatenation of unmasked latent $\bm{z}_{0}$.
-Moreover, the latent augmentation makes the image undergo two-round encoding and decoding from KL-VAE.
-This not only facilitates the incorporation of natural VAE degradation into the decoder training but also enhances the overall robustness of the model.
-We empirically apply 50% to operate the latent augmentation to training images.
-We fine-tune the decoder based on [64], showcasing superior consistency as compared in Figure 6.
+To balance the efficiency and efficacy, we design a one-step estimation of SD.
+As our target is to capture the generation gap, we use the clean latent $\bm{z}_{0}$ and all-zero mask $\mathbf{M}$ as conditions.
+This tells the SD all the needed information to generate the clean latent, ensuring the generated latent preserves content and only shift from the distribution gap.
+We follow the standard pipeline to estimate $\bm{z}_{0}$ with modified conditions as:
+$\hat{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\bm{z}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
+where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the prescribed variance schedule in SD; $\varepsilon_{\theta}(\cdot)$ is the frozen SD take as inputs noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$.
+The large step denoising is chosen to increase the distribution gap, as empirically SD could produce reliable results in small $t$ given the unmasked latent condition $\bm{z}_{0}$.
+Then we decode $\hat{\bm{z}}_{0}$ to image as the latent augmented inputs.
+This makes latent augmentation an off-line strategy.
+We apply latent augmentation to 50% training images.
+We fine-tune the decoder based on [62], showcasing superior consistency as compared in Fig.
+5(a).
+(a) w/ v.s. w/o latent aug.
+(b) Different image domains in MISATO.
+Figure 7: (a) Ablation of latent augmentation; (b) MISATO evaluation examples.
 4 Evaluation Dataset
-Figure 7: Examples of masked images constructed in MISATO. The first three cases focus on different image domains, while in the last case we focus on the inpainting of background.
-To validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [29], MegaDept[27], and COCO 2014 [30].
-Matterport3D contains a indoor image dataset, Flickr-Landscape is consisted of outdoor landscape images, MegaDepth involves a dataset of buildings, and COCO has precise segmentation of foreground subjects.
-For each dataset, we select 500 representative examples, yielding 2000 testing examples of size $512^{2}$.
+[…]
+In addition to the standard benchmark Places 2 [61], to validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29].
+The first three datasets contains images from indoor, outdoor landscape and building, respectively.
+COCO has object segmentation.
+We select 500 representative examples of size $512^{2}$ from each dataset.
+The MISATO dataset is available at https://github.com/Yikai-Wang/asuka-misato.
 The principle of constructing MISATO is to select the most representative and diverse examples.
-To this end, for Matterport3D, Flickr-Landscape and MegaDepth, we use CLIP visual model [37] to extract semantic visual features.
+To this end, for first three datasets, we use CLIP visual model [37] to extract semantic visual features.
 Then we use BisectingKMeans [44] to cluster each dataset into 500 clusters, and select the cluster centers as the evaluation data.
@@
 To this end, for each data we identify the foreground with provided segmentation and remove it from the generated masks, yielding a dataset specified for purely background inpainting.
-Combined together, MISATO contains four inpainting domains, indoor, outdoor landscape, building, and background.
-As shown in Figure 7, for each evaluation image, we adopt the similar masking strategy as in Section 3.1 while not using the rectangle and complement rectangle masks.
+Combined together, MISATO contains 2000 examples from four inpainting domains, indoor, outdoor landscape, building, and background.
+As shown in Fig.
+6(b), we adopt the masking strategy as in Sec.
+3.1 but excluding the rectangle and complement rectangle masks.
 The masking ratio is set as $[0.2,0.8]$.
@@
 We train ASUKA on Places2 [61].
-For the MAE [20] used in ASUKA, we fine-tune from the MAE used in Cao et al.
-[8] with a batch size of 1024.
-We use four self-attention blocks for the alignment module, training from scratch with AdamW [32] optimizer of learning rate 5e-2 for 23K steps of batch size 64 with the standard diffusion objective.
-We use warm-up strategy to set initial $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2k training steps and then freeze $p=10\%$.
+For the MAE [19] used in ASUKA, we train on images of size $256^{2}$, which is efficient and produce context-stable guidance for SD to generate high-resolution images.
+We fine-tune from the pre-trained MAE with a batch size of 1024.
+We use four self-attention blocks for the alignment module, training from scratch with AdamW [31] of learning rate 5e-2 with the standard diffusion objective.
+We set $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2K training steps and then freeze.
 We fine-tune the decoder based on [62] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay.
-We use ColorJitter to implement color augmentation, with the setting of brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
+We use ColorJitter for color augmentation, with brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
 Evaluation metrics.
-We compare inpainting algorithms using the following metrics: Peak Signal-to-Noise Ratio (PSNR) to compare the ratio of real data to introduced noise, Structural Similarity Index Measure (SSIM) [47] to identify the perceived change in structural information, Learned Perceptual Image Patch Similarity (LPIPS) [60] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [21] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [61] to measure the human-inspired linear separability.
-5.1 Comparison on MISATO
-Figure 8: Qualitative comparison on MISATO.
-We compare ASUKA with Co-Mod [61], MAT [26], LaMa [45], MAE-FAR [8], and SD [39] in landscape (first two rows), indoor (3rd and 4th rows), building (5th and 6th rows) and background inpaintings (last two rows).
-SD performs unconditional generation.
-SD-text utilizes text prompt of “background”.
-SD-prompt utilizes trained prompt for inpainting task using the same training setting of ASUKA.
-Inpainting algorithms usually suffer from fidelity issue, while SD-based algorithms suffer from stability issue and generate random elements.
-In contrast, ASUKA achieves stable and faithful inpainting.
+We use the Learned Perceptual Image Patch Similarity (LPIPS) [58] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [20] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [59] to measure the human-inspired linear separability.
+Competitors.
+We consider three variants of SD [39]: i) SD, prompted by "", performing unconditional generation; ii) SD-text, prompted by "background" as we have no captions in inpainting; iii) SD-prompt, prompted by learnable tokens which are trained using the same training pipeline of ASUKA.
+We also compare with state-of-the-art inpainting algorithms Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7], and SD-Repaint [32].
+Table 1: Quantitative comparison on MISATO and Places 2.
+Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7] and SD-Repaint [32] are state-of-the-art inpainting methods.
+SD [39] performs unconditional generation.
+SD-text uses “background” text prompt to guide generation.
+SD-prompt uses learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation.
+ASUKA and SD variants use the stable diffusion text-guided inpainting model v1.5.
+Dataset MISATO (2k images) Places 2 (36.5k images)
 […]
-Co-Mod [61] 21.805 0.787 0.179 17.421 0.243 0.109
-MAT [26] 21.553 0.787 0.176 17.261 0.255 0.122
-LaMa [45] 23.677 0.820 0.155 15.436 0.260 0.135
-MAE-FAR [8] 24.150 0.821 0.142 13.283 0.282 0.153
-SD [39] 22.454 0.796 0.168 12.812 0.345 0.211
-SD-text 22.394 0.793 0.164 12.603 0.337 0.207
-SD-prompt 22.381 0.798 0.160 12.517 0.331 0.204
-ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
-Table 1: Quantitative comparison on MISATO.
-Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8] are state-of-the-art inpainting algorithms.
-SD [39] performs unconditional generation.
-SD-text utilizes text prompts of “background” to perform text-guided generation.
-SD-prompt utilizes learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation.
-ASUKA and SD use the stable diffusion text-guided inpainting model v1.5.
-Competitors.
-We first compare ASUKA with SD [39].
-Specifically, we consider three variants of SD: i) SD, which generates inpainting results with no guidance, and performs unconditional generation; ii) SD-text, which generates inpainting results guided by the text prompt of “background”; iii) SD-prompt, where we replace the text prompt with learnable tokens and train the tokens using the same training pipeline of ASUKA.
-Beyond SD, we also compare ASUKA with state-of-the-art inpainting algorithms, including Co-Mod [61], MAT [26], LaMa [45], and MAE-FAR [8].
-Quantitative comparison results are reported in Table 1.
+Co-Mod [59] 0.179 17.421 0.243 0.109 0.267 5.794 0.274 0.096
+MAT [25] 0.176 17.261 0.255 0.122 0.202 3.765 0.348 0.195
+LaMa [45] 0.155 15.436 0.260 0.135 0.202 6.693 0.247 0.050
+MAE-FAR [7] 0.142 13.283 0.282 0.153 0.174 3.559 0.307 0.105
+SD [39] 0.168 12.812 0.345 0.211 0.193 1.514 0.375 0.207
+SD-text 0.164 12.603 0.337 0.207 0.191 1.506 0.373 0.202
+SD-prompt 0.160 12.517 0.331 0.204 0.189 1.477 0.390 0.234
+SD-Repaint [32] 0.227 27.861 0.016 0.007 0.251 12.466 0.217 0.045
+ASUKA 0.150 11.460 0.368 0.256 0.183 1.230 0.413 0.287
+5.1 Comparison on Benchmarks
+Quantitative comparison results are reported in Tab.
+1.
 ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS.
 Particularly, the U-IDS and P-IDS coincide with human preference [59], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin.
-The inconsistency results between PSNR, SSIM, LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net.
-On the other hand, the PSNR and SSIM is known to have bad correlation with human preferences, as several illustrative examples shown in the appendix.
-In summary, the results validate the contribution of ASUKA that yields stable and faithful inpainting through frozen SD.
-Qualitative comparison examples are shown in Figure 8.
-i) The state-of-the-art inpainting algorithms usually suffer from fidelity issues, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row.
+The inconsistency results between LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net.
+In summary, the results validate the contribution of ASUKA that yields context-stable and visual-consistent inpainting through frozen SD.
+This improvement is both validated in in-distribution dataset Places 2 and out-of-distribution dataset MISATO.
+Qualitative comparison examples are shown in Fig.
+8.
+i) The state-of-the-art inpainting algorithms usually suffer from unnatural generation, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row.
 LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks.
-ii) The three variants of SD usually suffer from the stability issue and generate random new elements in the masked region, in almost all the illustrated images.
+ii) The SD variants usually suffer from the context-stability issue and hallucinate unreasonable objects, in almost all the illustrated images.
 iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows.
-On the other hand, SD generates random elements in the masked region.
-iv) In contrast, ASUKA enjoys superior stable and faithful inpainting.
+SD hallucinates random elements in the masked region.
+iv) In contrast, ASUKA enjoys context-stable and visual-consistent inpainting.
 5.2 Further Analysis of ASUKA
-[…]
-p.t. MAE 22.485 0.799 0.151 11.513 0.354 0.258
-f.t. MAE 22.495 0.799 0.150 11.460 0.368 0.256
-Table 2: Quantitative comparison of ASUKA using pre-trained MAE (p.t. MAE) from [8] and fine-tuned MAE (f.t. MAE).
 Ablation of MAE prior.
-We compare our fine-tuned MAE with directly adopting the MAE trained in [8] without further fine-tuning.
-To this end, we train ASUKA with the MAE in [8] using the same training strategy and compare the results in.
+We compare our fine-tuned MAE with directly adopting the MAE trained in [7].
+To this end, we train ASUKA with the MAE in [7] using the same training strategy and compare the results in Tab.
+2(a).
 Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS.
 This improvement comes from the better adaptation on the real-world masks.
-[…]
-linear 22.402 0.797 0.155 11.934 0.361 0.227
-attn 22.474 0.798 0.152 11.613 0.362 0.234
-cross x4 22.488 0.799 0.152 11.762 0.368 0.238
-ASUKA 22.495 0.799 0.150 11.460 0.368 0.256
-Table 3: Ablation of different alignment modules.
-Linear adopts linear layer to align feature dimension only; attn adopts a self-attention layer to further align distribution; cross x4 adopts 4 cross-attention layers; while ASUKA adopts 4 self-attention layers.
 Ablation of alignment module.
@@
 ASUKA adopts 4 self-attention blocks.
-Results are shown in Table 3.
-The self-attention block shows improved results on simply aligning the dimension as well as using a cross-attention block to align the distribution.
+Results are shown in Tab.
+2(b).
+The self-attention block shows improved results compared with only align dimension and cross-attention block.
 Using 4 self-attention blocks improves the capacity of the alignment module.
+Ablation of decoder.
+For the decoder, we compare ASUKA with i) VAE: the decoder used in SD; ii) + cond.: the decoder conditioned on unmasked image [62]; iii) + color: only use color augmentation to train the decoder; Results are in Tab.
+2(d), showing the superiority of our augmented decoder.
+Table 2: Ablation stuides of ASUKA.
+(a) Comparison of ASUKA using pre-trained (p.t.) MAE v.s. fine-tuned (f.t.) MAE.
 […]
-KL-VAE [39] 22.454 0.795 0.156 11.949 0.343 0.208
-Asy. VAE [64] 22.566 0.801 0.151 11.634 0.361 0.231
-color-aug 22.475 0.799 0.152 11.603 0.357 0.229
-Ours 22.495 0.799 0.150 11.460 0.368 0.256
-Table 4: Quantitative comparison of different decoders for SD.
-Ablation of decoder training pipeline.
-For the decoder, we compare ASUKA with i) KL-VAE: the decoder used in SD; ii) Asy.
-VAE: the decoder trained by [64]; iii) color-aug: only use color augmentation to train the decoder; Results are shown in Table 4, which shows the superiority of our augmented decoder to ensure the mask-unmask consistency.
+p.t. 0.151 11.513 0.354 0.258
+f.t. 0.150 11.460 0.368 0.256
+(b) Ablation of different alignment modules. Linear adopts linear layer; attn adopts a self-attention layer; cross x4 adopts 4 cross-attention layers; ASUKA adopts 4 self-attention layers.
+[…]
+linear 0.155 11.934 0.361 0.227
+attn 0.152 11.613 0.362 0.234
+cross x4 0.152 11.762 0.368 0.238
+ASUKA 0.150 11.460 0.368 0.256
+(c) User-study of top-1 ratio among all the inpainting results.
+Context-stability (C.S.) measures the coherence between masked region and unmasked surroundings, with a preference of not generating new elements; Visual-consistency (V.C.) measures the color consistency.
+Model C.S.(%) V.C.(%)
+Co-Mod [59] 3.98 4.98
+MAT [25] 7.40 3.20
+LaMa [45] 8.18 8.28
+MAE-FAR [7] 4.88 5.60
+SD [39] 10.58 5.75
+SD-text 7.70 15.83
+SD-prompt 16.18 15.78
+SD-Repaint [32] 1.60 0.55
+ASUKA 39.43 40.05
+(d) Comparison of different decoders for SD.
+VAE [39] is the original decoder used by SD; + cond.
+[62] is the decoder conditioned on unmasked image; + color uses the color augmentation strategy to perform local harmonization task; Ours further combines latent augmentation strategy to handle the gap between generated latent and real latent.
+[…]
+VAE 0.156 11.949 0.343 0.208
++ cond. 0.151 11.634 0.361 0.231
++ color 0.152 11.603 0.357 0.229
+Ours 0.150 11.460 0.368 0.256
+User-Study.
+To evaluate the user preference on inpainting algorithms, we conduct an user-study.
+Specifically, we randomly select 40 testing images.
+We ask the user to select the best inpainting results from the following perspectives respectively: i) Context-stability (C.S.): the generated region should be context-stable with surrounding unmasked region, with a preference of not generating new elements; ii) Visual-consistency (V.C.) : the visual consistency between masked and unmasked regions.
+We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms in Tab.
+2(c).
+This result validate the efficacy of ASUKA on alignment with human preference.
 Figure 9: The curse of self-attention, causing the MAE falsely estimate the masked region and powerful text-guided diffusion models fail to generation content based on text prompts.
 ASUKA potential circumvents this issue by using a blank paper image as the input to the MAE to provide correct prior.
-Limitation: The “curse” of self-attention.
+Limitation: The "curse" of self-attention.
 The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module.
-Specifically, as showin in Figure 9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal.
+Specifically, as shown in Fig.
+9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal.
 Notably, this curse of self-attention significantly impacts diffusion-based generative models.
-It results in the inability to accurately follow “blank paper” text prompts, even when employing a substantial classifier-free guidance scale of 9.
+It results in the inability to accurately follow "blank paper" text prompts, even when employing a substantial classifier-free guidance scale of 9.
 This issue is not unique to SD but is also a common problem in other advanced text-guided diffusion models, such as OpenAI’s DALL-E 2 [38] and Adobe’s FileFly [1].
@@
 6 Conclusion
-In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve stable and faithful inpainting via frozen Stable Diffusion inpainting model (SD).
-To achieve stability, we adopt a reconstruction-based masked-auto encoder (MAE) as the stable prior for masked region purely from unmasked region.
-Then we align the stable prior to frozen SD with the proposed alignment module.
-To achieve fidelity, we resolve the mask-unmask inconsistency of SD in the decoding process.
-We fine-tune a unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process.
+In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve context-stable and visual-consistent inpainting via frozen Stable Diffusion inpainting model (SD).
+To achieve context-stability, we adopt a reconstruction-based masked auto-encoder (MAE) as the context-stable prior for masked region purely from unmasked region.
+Then we align the context-stable prior to frozen SD with the proposed alignment module.
+To achieve visual-consistency, we resolve the mask-unmask color inconsistency of SD in the decoding process.
+We train an unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process.
 To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets.
-ASUKA enjoys superior stable and faithful inpainting results than state-of-the-art inpainting algorithms and SD-based variants.
-Supplementary Material
-7 Extra Analysis
-7.1 Comparison on Places2
-[…]
-Co-Mod [61] 19.413 0.709 0.267 5.794 0.274 0.096
-MAT [26] 19.729 0.740 0.202 3.765 0.348 0.195
-LaMa [45] 21.503 0.771 0.202 6.693 0.247 0.050
-MAE-FAR [8] 22.140 0.779 0.174 3.559 0.307 0.105
-SD [39] 20.302 0.746 0.193 1.514 0.375 0.207
-SD-text 20.359 0.748 0.191 1.506 0.373 0.202
-SD-prompt 20.548 0.753 0.189 1.477 0.390 0.234
-ASUKA 20.406 0.752 0.183 1.230 0.413 0.287
-Table 5: Quantitative comparison on Places 2.
-We evaluation ASUKA and other inpainting algorithms on the standard testing benchmark dataset of Places2 [63], which includes 36500 testing examples.
-Results on this dataset also validate the efficacy of ASUKA.
-7.2 User-Study
-Model Consistency (%) Quality (%)
-Co-Mod [61] 4.69 5
-MAT [26] 7.5 3.33
-LaMa [45] 8.44 8.33
-MAE-FAR [8] 5.31 5.83
-SD [39] 10.63 5.83
-SD-text 7.81 15.83
-SD-prompt 16.25 15.83
-ASUKA 39.38 40
-Table 6: User-study of top-1 ratio on consistency and quality among all the inpainting results.
-To evaluate the user preference on inpainting algorithms, we conduct an user-study.
-Specifically, we randomly select 40 testing images along with inpainting results on these images.
-We ask the user to select the best inpainting results from the following perspectives respectively: i) Consistency: the masked region should be consistent with unmasked region, involving color and texture consistency and not generating new elements; ii) Quality: the fidelity of the inpainting results.
-The first metric measures the stability while the other measures the fidelity of the inpainting results.
-We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms.
-This result validate the efficacy of ASUKA on alignment with human preference.
-7.3 The Misalignment Between PSNR, SSIM and Human Preference
-The PSNR and SSIM are known to be have bad correlation with human preferences.
-To show this, we report the PSNR and SSIM scores for the qualitative comparison on MISATO ( Figure 8 in the paper) in Figure 10.
-Higher PSNR and SSIM do not always indicate better results that human prefers.
-Particularly, blurred inpainting results usually have higher PSNR and SSIM scores, like the 1st, 2rd, 4th, 8th rows.
-However, it is clear that blurred results are not expected.
-Figure 10: Misalignment between PSNR, SSIM and human preference. Images are same with Figure 8. PSNR and SSIM scores are listed in the bottom, where the best values are marked as red.
+ASUKA enjoys context-stable and visual-consistent inpainting results and superior than state-of-the-art inpainting algorithms and SD-based variants.

### v3

@@
-Project Page: https://yikai-wang.github.io/asuka/ MISATO Dataset: https://github.com/Yikai-Wang/asuka-misato
+Towards Enhanced Image Inpainting: Mitigating Unwanted Object Insertion and Preserving Color Consistency
 Abstract
-Recent progress in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing large irregular masks.
-However, this enhanced generation often introduces context-instability, leading to arbitrary object generation within masked regions.
-This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generation capacity.
-Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a Masked Auto-Encoder (MAE) to produce reconstruction-based prior.
-Aligned with the powerful Stable Diffusion inpainting model (SD), ASUKA significantly improves context stability.
-ASUKA further adopts an inpainting-specialized decoder, highly reducing the color inconsistency issue of SD and thus ensuring more visual-consistent inpainting.
-We validate effectiveness of inpainting algorithms on benchmark dataset Places 2 and a collection of several existing datasets, dubbed MISATO, across diverse domains and masking scenarios.
-Results on these benchmark datasets confirm ASUKA’s efficacy in both context-stability and visual-consistency compared to SD and other inpainting algorithms.
-Keywords:
-Image Inpainting Context-Stability Visual-Consistency
-Figure 1: Comparison on $1024^{2}$ image (depicted from MegaDepth [26]) between our proposed ASUKA and reconstruction-based (MAE [19]) as well as generative-based (MAE-FAR [7], MAT [25], Co-Mod [59], LaMa [45], Stable Diffusion [39]) inpainting models.
-MAE offers stable estimates, yet falls short in texture detail.
-GAN-based inpainting struggles with low fidelity.
-Powerful SD hallucinates random elements and suffers from mask-unmask visual inconsistency (color of clouds in the orange square).
-ASUKA achieves context-stable and visual-consistent inpainting.
+Recent advances in image inpainting increasingly use generative models to handle large irregular masks.
+However, these models can create unrealistic inpainted images due to two main issues: (1) Unwanted object insertion: Even with unmasked areas as context, generative models may still generate arbitrary objects in the masked region that don’t align with the rest of the image.
+(2) Color inconsistency: Inpainted regions often have color shifts that causes a smeared appearance, reducing image quality.
+Retraining the generative model could help solve these issues, but it’s costly since state-of-the-art latent-based diffusion and rectified flow models require a three-stage training process: training a VAE, training a generative U-Net or transformer, and fine-tuning for inpainting.
+Instead, this paper proposes a post-processing approach, dubbed as ASUKA (Aligned Stable inpainting with UnKnown Areas prior), to improve inpainting models.
+To address unwanted object insertion, we leverage a Masked Auto-Encoder (MAE) for reconstruction-based priors.
+This mitigates object hallucination while maintaining the model’s generation capabilities.
+To address color inconsistency, we propose a specialized VAE decoder that treats latent-to-image decoding as a local harmonization task, significantly reducing color shifts for color-consistent inpainting.
+We validate ASUKA on SD 1.5 and FLUX inpainting variants with Places2 and MISATO, our proposed diverse collection of datasets.
+Results show that ASUKA mitigates object hallucination and improves color consistency over standard diffusion and rectified flow models and other inpainting methods.
+footnotetext: First three authors contribute equally. Most parts of this work was done when Yikai was at Fudan. Yanwei Fu is the corresponding author.
 1 Introduction
-Figure 2: Object removal on $1024^{2}$ images. ASUKA enjoys context-stable (first to second rows) and visual-consistent (second to third rows) inpainting compared with SD.
+Figure 1: Image inpainting on $1024^{2}$ images.
+ASUKA solves two issues existed in current diffusion and rectified flow inpainting models: (1) Unwanted object insertion, where randomly elements that are not aligned with the unmasked region are generated; (2) Color-inconsistency: the color shift of the generated masked region, causing smear-like traces.
+ASUKA proposes a post-training procedure for these models, significantly mitigates object hallucination and improves color consistency of inpainted results.
 Image inpainting [6] fills masked areas of images while maintaining consistency with the unmasked regions.
-Traditional inpainting algorithms [5, 13, 18, 24, 40] often result in blurred synthesis when reconstructing masked regions, as indicated by [36].
-The advent of Generative Adversarial Networks (GANs) has led to works [35, 27, 8, 45, 46, 59, 17, 21, 15] that utilize GANs to fill complex mask structures, achieving impressive inpainting results.
+Traditional inpainting algorithms [6, 21, 30, 41, 69] often result in blurred synthesis when reconstructing masked regions [62].
+The Generative Adversarial Networks (GANs) based models could fill complex mask structures, achieving impressive inpainting results [59, 46, 9, 75, 78, 99, 27, 33, 24].
 However, they still struggle with general challenging inpainting cases, particularly in filling large holes.
-The recent Stable Diffusion Inpainting (SD) [39] model, with its large model capacity and extensive dataset training, acts as a versatile tool for image inpainting.
-However, we have observed a phenomenon of "context-instability" problem, where the model generates random, unreasonable elements to fill masked regions of images, as depicted in first to second rows in Fig.
-2, as well as Fig.
-1 and Fig.
-8.
-This issue appears to come from the random masking strategy used to train SD.
-This strategy introduces training cases where foreground objects are completely masked and the SD is forced to fill masked regions with foreground objects.
-Consequently, SD will hallucinate unreasonable objects devoid of contextual information.
+Recently, more powerful generative model like Stable Diffusion [67] and FLUX [40], with their large model capacity and extensive training dataset, act as versatile tools for image inpainting.
+These models mainly follow the latent generation pipeline, first encode the image into a small latent space, then train the inpainting model in this latent space.
+However, these latent-based generative inpainting models still suffer from some issues, causing the inpainted image lacking fidelity. In particular:
+(1) The unwanted object insertion problem, where the model generates random, unreasonable elements to fill masked regions, as depicted in first to second rows in Fig.
+1.
+This issue comes from the random masking strategy used to train generative models.
+This strategy introduces training cases where foreground objects are completely masked but the models are forced to fill masked regions with foreground objects.
+Consequently, these models will hallucinate unreasonable objects devoid of contextual information.
 Adjusting prompts may reduce this risk, but the best prompt is image-dependent, making it infeasible for practical applications.
-Furthermore, the inpainted results of SD may suffer from "visual inconsistency" problem.
-This problem, less explored in academia but critical for real-world industrial applications, results in significant color discrepancies between inpainted and unmasked regions and exhibits smear-like traces in the image, as shown in the second to third rows in Fig.
-2, as well as Fig.
-1 and Fig.
-8.
-Essentially, this visual-inconsistency comes from the misalignment between the pixel distributions of filled results and original images due to information loss in the Kullback-Leibler Variational Autoencoder (KL-VAE) used in SD, as illustrated in Fig.
-5.
-To improve visual consistency, an inpainting-specialized VAE decoder is necessary to explicitly mitigate information loss based on available ground-truth information of unmasked images.
-To efficiently tackle context-instability and visual inconsistency issues, we enhance the SD [39] with regression-based reconstruction and distribution-aligned generation.
-This results in an improved image inpainting model that avoids generating unreasonable elements in the masked region and reduces mask-unmask color inconsistency.
-The vanilla SD [39] adopts a KL-VAE to compress image into latent and perform diffusion in the latent space.
-We manipulate its generation and decoding processes for better context stability and visual consistency.
-i) We propose using the Masked Auto-Encoder (MAE)[19] as a prior to guide and stabilize the SD’s generation process.
+(2) Furthermore, the inpainted results of latent inpainting models suffer from color inconsistency problem.
+This problem, less explored in academia but critical for real-world applications, results in color discrepancies between inpainted and unmasked regions, including mismatches of brightness, saturation, luminance and hue, and exhibits smear-like traces in the image, as shown in the second to third rows in Fig.
+1.
+Essentially, this color-inconsistency comes from the misalignment between the pixel distributions of filled results and original images due to imperfect latent generative model and VAE, as illustrated in Fig.
+4.
+Notably, this issue is not a big problem for generation, given that the whole image is generated, and the color shift is consistent across pixels.
+However, this issue is important for inpainting tasks, as we have ground-truth pixels for unmasked regions.
+When we replace the unmasked regions of the generated image with the ground-truth pixels, the color inconsistency largely influence the fidelity of the image.
+This issue may be solved by training a better VAE and explicitly enforce the color consistency.
+However, training or fine-tuning the VAE encoder introduces the subsequent fine-tuning of the latent generative models to match the new latent space, which is costly.
+In this paper, we propose to freeze the VAE encoder and the latent generative models, while fine-tuning the VAE decoder to improve color consistency.
+Specifically, we reformulate the decoding from latent to image as a local harmonization task, explicitly reduce the color inconsistency.
+Formally, to mitigate object hallucination and enhance the color-consistency of image inpainting models, we present the Aligned Stable inpainting with UnKnown Areas prior (ASUKA) framework.
+ASUKA enhances the latent inpainting models with regression-based reconstruction and distribution-aligned generation.
+This results in improved image inpainting models that avoid generating unreasonable elements in the masked region and reduces mask-unmask color inconsistency.
+The stable diffusion models [67] and the rectified flow models [40] adopt a VAE to compress image into latent and perform inpainting in the latent space.
+We manipulate their generation and decoding processes to reduce object hallucination and improve color consistency.
+We propose using the Masked Auto-Encoder (MAE)[31] as a prior to guide and stabilize the generation process.
 As shown in Fig.
-2, MAE yields stable yet blurred results, while SD may produce implausible content despite its impressive generation capacity.
-By aligning MAE prior with SD objectives, we improve context-stability without compromising performance.
-ii) We redesign the KL-VAE decoder in SD to address color inconsistencies between masked and unmasked regions.
-The new decoder, acting as a local harmonization model conditioned on unmasked image pixels, significantly enhances visual consistency.
-Formally, to enhance the context-stability and visual-consistency of image inpainting model, we present the Aligned Stable Inpainting with Unknown Areas prior (ASUKA) framework.
-i) To enhance context-stability, we replace the text-condition part of the SD with the MAE prior.
-We fine-tune the MAE to ensure adaptability to practical inpainting scenarios, as MAE is trained on uniformly distributed masks but inpainting usually contains large continuous masks.
-We train an alignment module to align MAE prior with SD using the standard SD objective while freezing the fine-tuned MAE and SD.
-ii) To address visual-consistency issues during decoding, we redesign the decoder of the KL-VAE in SD.
-We inject additional inputs of unmasked image pixels and 0-1 masks to the decoder.
-This new decoder is trained to perform local harmonization task, taking as inputs of inharmonious latent and taking as outputs the original harmonious image.
-We generate inharmonious latent via the proposed color and latent augmentation techniques to incorporate KL-VAE information loss and generation artifacts of SD.
-Then the decoder is trained to recover the original image from corrupted latent based on the ground-truth information of unmasked image.
-These steps collectively enable ASUKA to achieve more context-stable and visual-consistent inpainting results.
+1, MAE yields stable yet blurred results, while generative models may produce implausible content despite their impressive generation capacity.
+By aligning MAE prior with latent generative models, we reduce object hallucination without damaging performance.
+We redesign the VAE decoder to address color inconsistencies between masked and unmasked regions by acting as a local harmonization model conditioned on unmasked image pixels.
+Our decoder can be used as a plug-and-play module to improve general inpainting models, such as text-guided inpainting.
+These steps collectively enable ASUKA to achieve less object hallucination and more color-consistent inpainting results.
+We adopt ASUKA on two typical inpaitning models, Stable Diffusion v1.5 [67] and FLUX [40], to validate the generalization ability of ASUKA on different generation architectures.
 To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, in addition to the benchmark dataset Places 2 [101], we further utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [13], Flickr-Landscape [47], MegaDept[45], and COCO 2014 [48].
-This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation.
-Experiments on MISATO and Places 2, especially with large irregular masks, validate the efficacy of ASUKA.
-Contributions.
-ASUKA enhances image inpainting from context-stability and visual-consistency while leveraging the generation capacity of the frozen SD inpainting model.
-It achieves this through two main components: i) Context-Stable Alignment: ASUKA aligns the stable MAE prior with SD to provide a context-stable estimation of masked regions, replacing the text-condition part of SD with MAE prior.
-ii) Visual-Consistent Alignment: ASUKA re-formulates the decoding from SD latent to image as a local harmonization task, trains an inpainting-specialized decoder to align masked and unmasked regions during decoding and thus mitigates color inconsistencies caused by information loss from the SD.
+This dataset covers four distinct domains—landscape, indoor, building, and background—making it diverse to serve as a benchmark for evaluation.
+Experiments on MISATO and Places 2 with large irregular masks validate the efficacy of ASUKA.
+Contributions
+ASUKA enhances image inpainting with color-consistency and mitigate object hallucination while leveraging the generation capacity of the frozen inpainting model.
+It achieves this through two main components: (1) Context-Stable Alignment: ASUKA aligns the stable MAE prior with generative models to provide a context-stable estimation of masked regions, replacing the text-condition with MAE prior.
+(2) Color-Consistent Alignment: ASUKA re-formulates the decoding from latent to image as a local harmonization task, trains an inpainting-specialized decoder to align masked and unmasked regions during decoding and thus mitigates color inconsistencies.
 2 Related Works
-Image inpainting is a longstanding challenge to infill the missing regions with consistent pixels.
-Classical approaches, relying on patchmatch [14, 4, 56] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes.
-GAN-based [17] inpainting methods [36, 53, 59, 25, 7] have evolved with adaptive convolutions [30, 53, 54], attention mechanism [52, 51, 55, 23], and high-resolution generalization in frequency learning [45, 50, 12].
-Methods like Co-Mod [59] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [25, 60], yielding robust and high-quality generations, even for large masks.
-However, they suffer from unstable results and uncontrollable hallucinations conditioned on different random latent codes.
-On the other hand, methods emphasizing higher reconstruction penalties [35, 45, 7] achieve more stable generations, but struggle with large masks, often resulting in blur artifacts.
-Recent diffusion models [41, 39] have rapidly developed and demonstrated impressive inpainting results.
-However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction.
-As a consequence, diffusion-based approaches also suffer from context-instability issue.
-Adapting Stable Diffusion.
-SD [39] is one of the most widely used diffusion models.
-SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space.
-Various tuning and aligning methods have been developed to introduce additional conditions to SD, including image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [16], LoRA fine-tuning [22], and controlnet [57] to add other types of conditions.
-Our objective is to enhance context-stability while preserving the generation capacity of SD.
-Consequently, we refrain from fine-tuning the U-Net backbone of SD.
-The text-condition is not utilized in inpainting tasks.
-Textual-inversion [16] introduces learnable tokens in Clip text encoders to improve subject description.
-ObjectStitch [43] employs the CLIP visual feature to represent the subject.
-In our approach, we advocate removing the text-condition and instead guiding SD using the Masked Auto-Encoder (MAE) prior for masked region.
-Information loss in SD latent space.
-Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images.
+Image inpainting
+is the task of filling missing image regions with consistent pixels.
+Traditional methods using patch matching [22, 5, 95] or differential equations [6, 12, 7] focus on low-level features and often struggle with large gaps.
+GAN [27]-based inpainting [62, 91, 99, 44, 10] introduces adaptive convolutions [50, 91, 93], attention [90, 89, 92, 39], and frequency-based learning for high-resolution results [75, 87, 17].
+Methods like Co-Mod [99] address the challenging ill-posed inpainting issue [44, 100] and improve realism but may produce unstable outputs or unwanted artifacts due to random latent variables.
+Techniques with higher reconstruction penalties [59, 75, 10] offer more stability but can appear blurry on larger missing areas.
+Recent diffusion models [70, 67, 3] and rectified flow models [25, 40] achieve impressive results yet share GANs’ limitation of learning distributions over exact pixel alignment, which leads to unwanted object insertion.
+Adapting latent generative models
+Latent diffusion models (LDMs) [67] and rectified flow models [25, 40] are popular due to their ability to encode image semantics at lower resolutions by combining a VAE to learn a latent space and a generative model within this space.
+Various methods have been developed to introduce new conditions to these models, such as image-inversion for text-guided image translation [56], textual-inversion for personalization [26], LoRA fine-tuning [34], and controlnet [96] to add diverse conditions.
+Our goal is to mitgate object hallucination while preserving generation quality, so we avoid fine-tuning the generative backbone.
+For inpainting, we remove the text condition and instead guide the generation using a Masked Auto-Encoder [31] prior for masked regions.
+Information loss in latent inpainting models
+Although claimed only eliminates imperceptible details, the VAE used by diffusion and rectified flow models causes distortion in the reconstruction of images.
+In addition, the gap between generated latent and real latent also causes the color inconsistency.
 See Fig.
-5(a) for illustrative examples.
+4 for illustrative examples.
 OpenAI [61] proposes a larger decoder to improve the decoding quality of SD’s latent.
-Luo et al.
+Luo et.
+al.
 [55] propose a frequency-augmented decoder to address the super-resolution case.
-Zhu et al.
+Zhu et.
+al.
 [103] propose to preserve unmasked regions during decoding.
-In this paper, we ensure the low-frequency color visual-consistency in the decoding process of SD.
-Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning.
+In this paper, we ensure the low-frequency color color-consistency in the decoding process.
+Masked Image-Modeling [4]
+(MIM) is an active research area in self-supervised learning.
 Typical MIM methods [4, 31, 86, 14] split images into visible and masked patches, learning to estimate masked patches from visible patches.
@@
 While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy.
-Cao et al.
+Cao et.
+al.
 [10] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies.
-In contrast, this paper uses MAE prior to enhance the context-stability of SD.
+In contrast, this paper uses MAE prior to enhance the context-stability of diffusion and rectified flow models.
+Image harmonization
+aims to blend a foreground object with a background image while keeping the final result realistic and visually consistent [76].
+This task is often treated as an image translation problem [102, 19, 28, 20, 29, 60, 79, 51, 57, 66].
+Similarly, our work addresses color inconsistency issues in latent generative models.
+However, unlike standard image harmonization, where inconsistencies arise from combining images from different sources and thus different real image distributions, color inconsistencies in latent generative models stem from imperfections in the VAE and the generative model itself.
+Object insertion and removal
+are two opposite tasks in image inpainting.
+Object insertion focuses on adding foreground objects to the image using various methods, such as shape-guided masks [94, 85], text prompts [80, 85, 8, 16], learnable prompts [81, 104, 16], extra network modules [15, 35], or reference images of objects [71], etc.
+Some studies also explore completing partial objects using reference images [11] or learnable prompts [81].
+Object removal, on the other hand, aims to erase unwanted objects from an image.
+Common approaches include attention reweighting [42] and learnable prompts [81, 104].
+These techniques can help create new datasets [23].
+On the other hand, creating new datasets can also benefit these tasks [84].
+While most research focuses on designing better inpainting models, our work takes a different approach.
+We analyze a fundamental problem with latent generative models: they often introduce unwanted objects in the inpainting area.
+We also propose solutions to address this issue.
 3 Methodology
-Problem setup.
+Problem setup
 Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region.
@@
 In this paper, we focus on the standard inpainting task without utilizing other conditions.
-We adopt stable diffusion inpainting model [39] and focus on the general issues of inpainting models, i) context-instability issue: unstable and uncontrollable hallucinations, yielding random elements generated in the masked region; ii) visual-inconsistency issue: mask-unmask color inconsistency issue, yielding smear-like traces in the masked region.
-Figure 3: ASUKA tackles the context-instability issue by adopting the MAE to provide a stable prior for frozen SD to maintain the generation capacity while increasing context-stability.
-For the visual-inconsistency issue, ASUKA utilizes an inpainting-specialized decoder to achieve mask-unmask color consistency when decoding SD latent.
-(a) Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
-(b) Alignment module.
-Figure 4: (a) Use MAE as initial value v.s. as condition. (b) Alignment module.
-Overview.
+We focus on the general issues of inpainting models, (1) unwanted object insertion: unstable and uncontrollable hallucinations, yielding random elements generated in the masked region; (2) color-inconsistency: mask-unmask color inconsistency issue, yielding smear-like traces in the masked region.
+Figure 2: ASUKA tackles the unwanted object insertion issue by adopting the MAE to provide a stable prior for frozen latent generative models to maintain the generation capacity while mitigating object hallucination.
+For the color-inconsistency issue, ASUKA utilizes an inpainting-specialized decoder to achieve mask-unmask color consistency when decoding latent.
+We evaluate our proposed solution on two inpainting models: the Stable Diffusion v1.5 inpainting model (SD) [67] and the Control-Net fine-tuned FLUX inpainting model (FLUX) [2].
+We provide a brief introduction of these models in the appendix.
+We will demonstrate that our ASUKA effectively improves unwanted object mitigation and color consistency of these models.
+Overview
 The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Fig.
+2(a).
+ASUKA adopts the pre-trained latent inpainting models.
+Our target is to mitigate object hallucination and provide more color-consistent inpainting results while fully exploiting the generation capacity of frozen models.
+ASUKA includes (1) a context-stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with generative models and (2) a color-consistent alignment to align ground-truth unmasked region with generated masked region during decoding.
+To this end, we freeze the latent generative models, while replacing the text-condition part with our proposed MAE prior to mitigate object hallucination.
+To align the MAE prior to generative models, we introduce an alignment module, trained via the training objective of generative models.
+Additionally, to align masked and unmasked regions during decoding and resolve the information loss issue from VAE decoder and generative model which causes mask-unmask color inconsistency, we train an inpainting-specialized decoder to decode the latent back to the image space for seamless color-consistency.
+Combined together, ASUKA achieves less object hallucination and more color-consistent inpainting.
+3.1 Mitigate Object Hallucination via Stable Prior
+3.1.1 Masked Auto-Encoder Prior
+Context-stable prior
+While recent generative models rely on random noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly.
+Some inpainting models also utilize the reconstruction loss to reconstruct the masked region, but they also incorporate other types of losses like perceptual-loss [75] which implicitly reduces the stability.
+In contrast, MAE is known to provide a context-stable estimation of masked regions based purely on the unmasked regions.
+In this paper, we utilize MAE to produce the stable prior such that the improvement of inpainted result can be explicitly attributed to the improvement of mitigating object hallucination.
+MAE as context-stable prior
+As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a context-stable prior for generative models to not generate new concepts.
+However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the inpainting models to inpaint in image-to-image style, as in Fig.
 3.
-ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39].
-Our target is to provide more context-stable and visual-consistent inpainting results while fully exploiting the generation capacity of SD.
-ASUKA includes i) a context-stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a visual-consistent alignment to align ground-truth unmasked region with generated masked region during decoding.
-To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase context-stability.
-To align the MAE prior to SD, we introduce an alignment module, which is trained via the training objective of SD.
-Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask color inconsistency, we train an inpainting-specialized decoder to decode the latent back to the image space for seamless visual-consistency.
-Combined together, ASUKA achieves more context-stable and visual-consistent inpainting.
-3.1 Masked Auto-Encoder Prior
-MAE as context-stable prior.
-While recent generative models rely on random noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly.
-In contrast, MAE is known to provide a context-stable estimation of masked regions based purely on the unmasked regions.
-As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a context-stable prior for generative models to not generate new concepts.
-However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the SD to perform inpainting in image-to-image style, as illustrated in Fig.
-3(a).
-To this end, we adopt the MAE to provide prior to stabilizing SD.
-Training MAE.
+To this end, we adopt the MAE to provide prior to stabilizing diffusion models.
+Train MAE
 The original MAE is trained to estimate random masks uniformly distributed in the image, while inpainting task usually contains large continuous masks.
-Inspired by Cao et al.
+Inspired by Cao et.
+al.
 [10], we fine-tune the MAE to inpainting masks.
@@
 This benefits ASUKA to tackle the large hole inpainting task.
-3.2 Align MAE Prior with SD for Context-Stability
-Replace text-condition with MAE prior.
-SD is not trained on MAE priors.
-Fortunately, the text-condition part of SD is not utilized in inpainting, as we do not assume a text condition for inpainting task.
-Hence, we propose to replace the text-condition part of SD with our proposed MAE prior condition.
-However, as we do not fine-tune the SD, it cannot aligns well with the MAE prior.
-To achieve a proper guidance for SD, we introduce the alignment module to align MAE with SD in both dimension and distribution perspective, as shown in 3(b).
-Dimension alignment.
+Figure 3: Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
+3.1.2 Align MAE Prior with Generator
+Replace text-condition with MAE prior
+Generative inpainting models are not trained on MAE priors.
+As we do not assume a text condition for inpainting task, we propose to replace the text-condition of generative models with our proposed MAE prior condition.
+However, as we do not fine-tune the generative models, they cannot directly align well with the MAE prior.
+Hence, we introduce the alignment module to align MAE with generative models in both dimension and distribution perspective, as shown in Fig.
+2(b).
+Dimension alignment
 Particularly, the MAE prior $F_{\mathrm{MAE}}$ is of size $N_{m}\times M_{m}$, where $N_{m}$ is the sequence length and $M_{m}$ is the feature dimension.
-To align it with the SD condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
-Distribution alignment.
-After aligning the dimension, we use self-attention blocks to learn to better guiding SD, leading to the condition $C_{\mathrm{MAE}}$ for SD.
-We train our alignment module using the standard diffusion objective with the same masking strategy used to train the MAE, keeping other modules frozen.
-Misalignment.
-When training the alignment module with the set (input image, MAE prior, SD result), misalignment issues may arise.
-For example, if an object is completely masked, the MAE will predict the masked area with background, whereas the SD is trained to recreate the object.
+To align it with the diffusion or flow condition of size $N_{s}\times M_{s}$, we adopt a linear layer to map the feature dimension from $M_{m}$ to $M_{s}$ and set $N_{s}=N_{m}$ to preserve the local MAE prior.
+Distribution alignment
+After aligning the dimension, we use self-attention blocks to learn to better guiding generative models, leading to the condition $C_{\mathrm{MAE}}$.
+We train our alignment module using the standard generation objective with the same masking strategy used to train the MAE, keeping other modules frozen.
+Handle misalignment
+When training the alignment module with the set (input image, MAE prior, inpaint result), misalignment may arise.
+For example, if an object is completely masked, the MAE will predict the masked area with background, whereas the generative models are trained to recreate the object.
 This difference can lead the alignment module to mistakenly disregard the MAE prior.
-To address this, we improve the SD’s adherence to the MAE prior by substituting the MAE predicted prior with the MAE reconstructed prior at a probability of $p$.
+To address this, we improve the generative models’ adherence to the MAE prior by substituting the MAE predicted prior with the MAE reconstructed prior at a probability of $p$.
 The MAE reconstructed prior involves using MAE to recreate the image without masking any area, ensuring MAE has access to all information needed for reconstruction.
-This approach helps train the alignment module to better guide the SD.
-Figure 5: The illustration of the information loss of KL-VAE. (a) The hue of the image is shifted. (b) KL-VAE still suffers from non-ignorable shifts in low-frequency fields.
-(a) Qualitative SD1.5 inpainting results decoded by (b) vanilla decoder of SD [39], (c) larger decoder [62], (d) our decoder. Our decoder largely alleviate the mask-unmask color inconsistency.
-(b) Decoder trained by local harmonization task, enhancing mask-unmask consistency by reconstructing original image guided by the unmasked region from augments in color and latent spaces.
-Figure 6: ASUKA trains a specialized-decoder to solve the color-inconsistency issue from the information loss of KL-VAE and domain gap between generated and real latent.
-3.3 Mask-Unmask Align during Decoding for Visual-Consistency
-Information loss of KL-VAE.
-SD performs all its generative processes in the latent space and subsequently decodes these latent codes back to image space using KL-VAE.
+This approach helps train the alignment module to better guidance.
+3.2 Enhancing Color-Consistency in Decoding
+3.2.1 Color-Inconsistency
+Color-inconsistency is a general problem
+The color-inconsistency between masked and unmasked regions is a general problem in generative inpainting models.
+This inconsistency comes when the generative masked region suffers from a color shift compared with the unmasked region.
+As in Fig.
+4, the color shift happens in all kinds of scenarios, including indoor and outdoor scenes, random or continuous masks, and may cause darker or lighter color shift.
+This shift comes from the imperfect VAE and latent generator.
+Information loss of VAE
+Popular latent diffusion and rectified flow models perform all the generative processes in the latent space and subsequently decodes these latent codes back to image space using VAE.
 Despite the decoder being trained to reconstruct the image, it encounters challenges associated with information loss.
 Particularly in tasks like inpainting, we have ground-truth values for the unmasked region.
-Though Rombach et al.
-[39] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE and GAN are used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in KL-VAE could not be neglected, as verified in Fig.
+Though Rombach et.
+al.
+[67] claimed that the diffusion model should prioritize the informative semantic compression, while the VAE is used to tackle perceptual compression with high-frequency details, we argue that low-frequency semantic loss in VAE could not be neglected, as verified in Fig.
 6 (b).
-The KL-VAE will not only noticeably degrades high-frequency details but also shifts in semantic hue.
-This shift can be verified by repeated reconstruction with KL-VAE, as shown in Fig.
+The VAE will not only noticeably degrades high-frequency details but also shifts in colors.
+This shift can be verified by repeated reconstruction with VAE, as shown in Fig.
 6 (a) where larger shift is observed during repeated reconstruction.
-As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant visual-inconsistencies.
-This visual-inconsistency issue is more severe under the irregular or large mask structures.
-Unmask-region conditioned decoder.
-The aforementioned information loss is prohibitive to be addressed by the SD U-Net alone, especially given the already degraded input latent codes.
-Zhu et al.
-[62] adopts a larger decoder, with additional inputs of masked images.
+As human is sensitive to low-frequency information changes in the image, even subtle color shifts can induce significant inconsistencies.
+This issue is more severe in irregular or large mask cases.
+Figure 4: The color shift exists in all kinds of scenarios in inpainted images, including indoor and outdoor scenes, random or continuous masks, and may cause darker or lighter color shift.
+Figure 5: Inpainting w/ v.s. w/o latent augmentation. The latent augmentation handles the gap between generated and real latent.
+Gap between real and generated latents
+Apart from the information loss of VAE in reconstruction, there is another gap between the generated and real latents.
+This gap also causes color inconsistency even if we alleviate the VAE reconstruction loss, as in Fig.
+5.
+We need to solve both the loss of VAE and the latent generator for better color-consistency.
+Figure 6: (a) The color of the reconstruted image is shifted, where larger shift is observed during repeated reconstruction. (b) VAE suffers from non-ignorable shifts in low-frequency fields.
+Figure 7: Decoder trained by local harmonization task, enhancing mask-unmask consistency by reconstructing original image guided by the unmasked region from augments in color and latent spaces.
+Figure 8: SD1.5 inpainting results decoded by (b) vanilla decoder of SD [67], (c) conditional decoder [103], (d) our decoder. Our decoder largely alleviate the mask-unmask color inconsistency.
+3.2.2 Mask-Unmask Align during Decoding
+We propose to solve the color-inconsistency and ensure the mask-unmask alignment during VAE decoding.
+Unmask-region conditioned decoder
+The basic solution is to incorporate the ground-truth unmasked region in the decoding, then we could have access to the unbiased color information.
+Zhu et.
+al.
+[103] adopts decoder with additional inputs of masked images.
 However, it still fails to handle the incompatible color and texture between the original images and compressed ones in challenging scenes as verified in Fig.
-5(a) (c).
-The low-frequency gap between degraded and original images makes it challenging to explicitly address this issue.
-To solve the mask-unmask visual-inconsistency issue, we reformulate the decoding as a local harmonization task to preserve the visual consistency.
-As shown in Fig.
-3, compared with SD decoder, our decoder involves additional inputs of masked images in the pixel-wise color space and the 0-1 mask.
-Mask-unmask visual-consistent decoder.
-To train the decoder to ensure visual-consistency between generated latent and unmasked pixels, we re-formulate the decoding as a local harmonization task.
-Specifically, we propose the color and latent augmentation as shown in Fig.
-5(b) to estimate and enlarge the visual-inconsistency during decoding.
-We follow the standard KL-VAE training pipeline of SD, but replacing the inputs with augmented ones.
-Particularly, we use the original image as the reconstruction target and use color and latent augmentation to corrupt input image, simulating the information loss of KL-VAE and domain gap between generated and real latent, respectively.
-This forces the decoder to reconstruct the un-corrupted image based on the ground-truth unmasked region.
-i) Color augmentation: The first information loss comes from the imperfect VAE reconstruction.
-VAE is trained to encode and decode the input image for reconstruction, but suffers from color shift as shown in Fig.
-5(a) (b).
+8 (c).
+The gap between degraded and original images makes it challenging to explicitly address this issue.
+Mask-unmask color-consistent decoder
+To train the decoder to ensure color-consistency between generated latent and unmasked pixels, we re-formulate the decoding as a local harmonization task.
+Our decoder involves additional inputs of masked images in the pixel-wise color space and the 0-1 mask.
+To properly train the decoder, we propose the color and latent augmentation as shown in Fig.
+7 to estimate and enlarge the color-inconsistency.
+We follow the standard VAE training pipeline, but replacing the inputs with augmented ones.
+Particularly, we use the original image as the reconstruction target and use color and latent augmentation to corrupt input image, simulating the information loss of VAE and domain gap between generated and real latent, respectively.
+This forces the decoder to reconstruct the clean image based on the ground-truth unmasked region.
+Color augmentation
+We use color augmentation to capture the VAE loss as in Fig.
+8 (b).
 Empirically, further conditioned on unmasked image alleviate but not solve the color inconsistency issue, as shown in Fig.
-5(a) (c).
+8 (c).
 Hence, we need to explicitly train the decoder to ensure color consistency.
 To this end, we augment all training images in brightness, contrast, saturation, and hue, and requires the decoder to reconstruct original image conditioned on the unaugmented unmasked image.
-This encourages the decoder to fully utilize the information from unmasked regions.
-ii) Latent augmentation: The other information loss comes from the distribution gap between SD generated latent and real latent.
-This gap also causes color inconsistency even if we alleviate the VAE reconstruction loss, as shown in Fig.
-6(a).
-To simulate this loss, we incorporate the artifacts generated from the SD to train the decoder.
+This encourages the decoder to faithfully follow the unmasked regions.
+Latent augmentation
+To simulate the gap between generated and real latent, we incorporate the artifacts generated from the generative models to train the decoder.
 However, denoising to real images iteratively is notably time-consuming, even with DDIM [73].
-To balance the efficiency and efficacy, we design a one-step estimation of SD.
+To balance the efficiency and efficacy, we design a one-step estimation.
 As our target is to capture the generation gap, we use the clean latent $\bm{z}_{0}$ and all-zero mask $\mathbf{M}$ as conditions.
-This tells the SD all the needed information to generate the clean latent, ensuring the generated latent preserves content and only shift from the distribution gap.
+This tells the generator all the needed information to generate the clean latent, ensuring the generated latent preserves content and only shift from the generation gap.
 We follow the standard pipeline to estimate $\bm{z}_{0}$ with modified conditions as:
-$\hat{\bm{z}}_{0}=\frac{1}{\sqrt{\overline{\alpha}_{t}}}(\bm{z}_{t}-\sqrt{1-\overline{\alpha}_{t}}\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
-where the timestep $t$ is randomly sampled from $[500,1000)$; $\overline{\alpha}_{t}$ indicates the prescribed variance schedule in SD; $\varepsilon_{\theta}(\cdot)$ is the frozen SD take as inputs noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$.
-The large step denoising is chosen to increase the distribution gap, as empirically SD could produce reliable results in small $t$ given the unmasked latent condition $\bm{z}_{0}$.
+$\hat{\bm{z}}_{0}=\frac{1}{a}(\bm{z}_{t}-b\varepsilon_{\theta}\left([\bm{z}_{t};\bm{z}_{0};\mathbf{M}],t\right)),$ (1)
+where the timestep $t$ is randomly sampled from $[500,1000)$; $a$ indicates the prescribed variance schedule, $a^{2}+b^{2}=1$ in diffusion models while $a+b=1$ in rectified flow models; $\varepsilon_{\theta}(\cdot)$ is the frozen generator take as inputs noised $\bm{z}_{t}$, unmasked $\bm{z}_{0}$, and all-zero masking $\mathbf{M}$.
+The large step denoising is chosen to increase the distribution gap, as empirically the generator could produce reliable results in small $t$ given the unmasked latent condition $\bm{z}_{0}$.
 Then we decode $\hat{\bm{z}}_{0}$ to image as the latent augmented inputs.
@@
 We apply latent augmentation to 50% training images.
-We fine-tune the decoder based on [62], showcasing superior consistency as compared in Fig.
-5(a).
-(a) w/ v.s. w/o latent aug.
-(b) Different image domains in MISATO.
-Figure 7: (a) Ablation of latent augmentation; (b) MISATO evaluation examples.
-4 Evaluation Dataset
+The fine-tuned decoder showcases superior consistency as compared in Fig.
+8.
+Figure 9: Inpainting results for 5122 images.
+GANs generate blurred results; SD variants hallucinate unreasonable objects and suffer from color shift.
+ASUKA achieves unwanted-object-mitigated and color-consistent inpainting.
+More results are in the appendix.
+4 Experiments
+Evaluation datasets
+We follow previous works to evaluate on the standard benchmark Places 2 [101] validation set of 36,500 images.
+In addition, to validate across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [13], Flickr-Landscape [47], MegaDept[45], and COCO 2014 [48] to handle indoor, outdoor, building and background inpainting, respectively.
+We select 500 representative examples of size $512^{2}$ and $1024^{2}$ from each dataset, forming a total of 2,000 testing examples.
+See details in the appendix.
+General evaluation metrics
+We use the Learned Perceptual Image Patch Similarity (LPIPS) [97] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [32] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [99] to measure the human-inspired linear separability.
+Evaluate object hallucination and color-consistency
+We introduce two new metrics to assess the object hallucination and color-consistency of inpainted images.
+(1) CLIP@mask (C@m): We use CLIP to get visual features from both the ground-truth and the inpainted masked region, then calculate their cosine similarity.
+Following the standard CLIP score, we multiply the result by 100 and clip negative values, yielding a range from 0 to 100.
+(2) Gradient@edge (G@e): We calculate the average pixel gradient difference along the edges of the masked region with respect to the ground-truth image to assess color smoothness.
+A smaller gradient difference means more similar color transitions to the ground-truth image and, therefore, less color shift.
+Table 1: Quantitative comparison on MISATO and Places 2. Top-3 results are colored.
+Dataset MISATO Places 2
 […]
-In addition to the standard benchmark Places 2 [61], to validate the inpainting performance across different domains and mask styles, we construct a evaluation dataset, dubbed as MISATO, from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29].
-The first three datasets contains images from indoor, outdoor landscape and building, respectively.
-COCO has object segmentation.
-We select 500 representative examples of size $512^{2}$ from each dataset.
-The MISATO dataset is available at https://github.com/Yikai-Wang/asuka-misato.
+Co-Mod [99] 0.179 17.421 0.243 0.109 0.924 52.106 0.267 5.794 0.274 0.096 0.951 166.914
+MAT [44] 0.176 17.261 0.255 0.122 0.925 48.722 0.202 3.765 0.348 0.195 0.955 163.442
+LaMa [75] 0.155 15.436 0.260 0.135 0.928 46.270 0.202 6.693 0.247 0.050 0.953 153.653
+MAE-FAR [10] 0.142 13.283 0.282 0.153 0.940 43.613 0.174 3.559 0.307 0.105 0.958 149.843
+SD-Repaint [54] 0.227 27.861 0.016 0.007 0.915 80.410 0.251 12.466 0.217 0.045 0.947 176.421
+SD [67] 0.168 12.812 0.345 0.211 0.951 63.844 0.193 1.514 0.375 0.207 0.959 160.705
+SD-text 0.164 12.603 0.337 0.207 0.952 63.776 0.191 1.506 0.373 0.202 0.959 160.418
+SD-token [81] 0.160 12.517 0.331 0.204 0.955 61.700 0.189 1.477 0.390 0.234 0.960 158.924
+SD-IP [88] 0.157 12.204 0.398 0.242 0.956 62.704 0.186 1.539 0.389 0.173 0.953 148.571
+SD-T2I [58] 0.166 13.806 0.365 0.222 0.949 63.866 0.195 1.720 0.384 0.160 0.951 148.549
+SD-CAEv2 [98] 0.157 29.179 0.193 0.045 0.901 69.890 0.192 6.887 0.287 0.065 0.921 151.863
+SD-LaMa [75] 0.157 12.159 0.390 0.256 0.956 62.726 0.188 1.522 0.389 0.168 0.953 148.461
+ASUKA-SD 0.150 11.495 0.423 0.312 0.958 47.753 0.183 1.230 0.413 0.287 0.961 147.733
+Competitors
+We primarily use the SD v1.5 inpainting model [67] to analyze and compare ASUKA with competitors, while validating ASUKA’s generalization ability with FLUX.
+We consider three SD v1.5 inpainting variants: SD: uses a null-prompt for unconditional generation; SD-text: uses "background" as a prompt since no captions are used in inpainting; SD-token [81]: uses learnable tokens trained with ASUKA’s pipeline.
+To test other ways of incorporating the MAE condition, we implement the following: SD-IP, uses IP-Adapter [88]; SD-T2I, uses T2I-Adapter [58]; SD-CAEv2, uses a CLIP-style alignment module CAEv2 [98]; We also test SD-LaMa, which inputs LaMa [75] inpainting results instead of MAE.
+We also compare with leading inpainting algorithms Co-Mod [99], MAT [44], LaMa [75], MAE-FAR [10], and SD-Repaint [54].
+We provide implementation details in the appendix.
+4.1 Comparison on Benchmarks
+Quantitative comparison
+Results on SD are reported in Tab.
+1.
+(1): Although ASUKA-SD is based on a fixed SD model, it consistently outperforms SD across all evaluation metrics, achieving state-of-the-art results in FID, U-IDS, and P-IDS.
+Notably, U-IDS and P-IDS are closely aligned with human preferences [99] and have a potential maximum score of 0.5, highlighting ASUKA’s strong performance.
+(2): Compared to other adapters that align the MAE prior with SD, ASUKA-SD shows consistently superior performance across all metrics.
+This demonstrates the effectiveness of our straightforward alignment module.
+(3): While the LaMa condition improves inpainting quality, as shown by FID and IDS scores, it is less effective than the MAE condition.
+When using the MAE condition as a prior, improvements can be attributed to better mitigation of object hallucination.
+(4): ASUKA-SD consistently performs better than all competitors on CLIP@mask, showcasing the strength of its improved mitigation of object hallucination.
+(5): Pixel-based GAN inpainting models perform better in the Gradient@edge metric, suggesting that color shifts may originate from the compressed latent space.
+ASUKA-SD, however, still shows significant improvements over all SD variants, highlighting its enhanced color consistency.
+(6): The second-to-best LPIPS scores are partially due to using a frozen SD, where ASUKA achieves consistent improvements but remains constrained by the frozen U-Net.
+These results confirm that ASUKA-SD improves color consistency and mitigation of object hallucination in inpainting, even when using frozen latent inpainting models.
+This advantage is evident both in the in-distribution dataset Places2 and the out-of-distribution dataset MISATO.
+Qualitative comparison
+examples are shown in Fig.
+9.
+(1) The state-of-the-art inpainting algorithms usually suffer from unnatural generation, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row.
+LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks.
+(2) The SD variants usually suffer from the unwanted object insertion issue and hallucinate unreasonable objects, in almost all the illustrated images.
+(3) In contrast, ASUKA enjoys unwanted-object-mitigated and color-consistent inpainting.
+4.2 Further Analysis of ASUKA
+In this part, we conduct more experiments to analysis ASUKA. More analysis can be found in the appendix.
+Extension to FLUX
+To demonstrate ASUKA’s versatility, we trained it on FLUX (see Tab. 2). ASUKA-FLUX consistently outperforms the original FLUX. Results on Places 2 and qualitative comparisons are in the appendix.
+Table 2: FLUX and ASUKA-FLUX on MISATO@512. Results on 1K and qualtitative results are in the appendix.
+[…]
+FLUX 0.254 12.839 0.351 0.223 0.951 65.928
+ASUKA-FLUX 0.206 11.372 0.428 0.327 0.962 48.635
+Table 3: Comparison of different decoders for SD.
+[…]
+VAE 0.156 11.949 0.387 0.253 0.953 63.142
++ cond. 0.151 11.634 0.410 0.272 0.955 48.588
++ color 0.152 11.603 0.407 0.273 0.954 49.538
+Ours 0.150 11.495 0.423 0.312 0.958 47.753
+Ablation of decoder
+For the decoder, we compare ASUKA-SD with (1) VAE: the decoder used in SD; (2) + cond.: the decoder conditioned on unmasked image [103]; (3) + color: only trained with color augmentation ; Results are in Tab.
+3, showing the superiority of our decoder.
+Table 4: Ablation of different alignment modules.
+[…]
+linear 0.155 11.934 0.400 0.263 0.953 48.983
+attn 0.152 11.613 0.403 0.268 0.954 48.785
+cross x4 0.152 11.762 0.405 0.256 0.953 48.279
+Ours 0.150 11.495 0.423 0.312 0.958 47.753
+Ablation of alignment module
+We validate the efficacy of our alignment module step by step: (1) linear: Use linear layer to align feature dimension only; (2) attn: Based on linear, further use a single self-attention block to align the distribution; (3) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior.
+ASUKA-SD adopts 4 self-attention blocks.
+Results are shown in Tab.
+4.
+The self-attention block shows improved results compared with only align dimension and cross-attention block.
+Using 4 self-attention blocks improves the capacity.
+5 Conclusion
+In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve unwanted-object-mitigated and color-consistent inpainting via frozen latent inpainting models.
+To avoid unwanted object insertion, we adopt a reconstruction-based masked auto-encoder (MAE) as the context-stable prior for masked region purely from unmasked region.
+Then we align the context-stable prior to frozen generative models with the proposed alignment module.
+To achieve color-consistency, we resolve the mask-unmask color inconsistency in the latent decoding process.
+We train an unmask-region conditioned VAE decoder to perform local harmonization during the decoding process.
+To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets.
+We propose two new metrics to explicitly evaluate the object hallucination and color-consistency of inpainted images.
+ASUKA enjoys unwanted-object-mitigated and color-consistent inpainting results and superior than leading inpainting models.
+Acknowledgments
+The authors would like to thank Huawei Ascend Cloud Ecological Development Project for the support of Ascend 910 processors.
+Supplementary Material
+6 Brief Introduction of Backbone Models
+We evaluate our proposed solution on two inpainting models: the Stable Diffusion v1.5 inpainting model (SD) [67] and the Control-Net fine-tuned FLUX inpainting model (FLUX) [2].
+Both models are representative latent inpainting models that use a VAE [38] to compress images into a smaller latent space.
+In SD, a diffusion process [72] maps the latent space to random Gaussian noise, and a U-Net [68] learns the reverse denoising path.
+Text condition is introduced through cross-attention layers [77].
+The inpainting version of SD extends the U-Net input by concatenating the masked image and mask with the noise along the channel dimension.
+Conversely, FLUX uses rectified flow [52, 1, 49] to map the latent space to noise and a vision transformer [63] for generation.
+Text condition is applied by concatenating text with image patches as transformer input, while a pooled text condition is injected into the normalization layers.
+Since the original FLUX [40] does not support inpainting, we use a Control-Net [96] fine-tuned version [2] that modifies FLUX’s transformer output by conditioning on the masked image and mask.
+We demonstrate that our ASUKA effectively improves unwanted object mitigation and color consistency of these models.
+7 Details about MISATO
 The principle of constructing MISATO is to select the most representative and diverse examples.
@@
 To this end, for each data we identify the foreground with provided segmentation and remove it from the generated masks, yielding a dataset specified for purely background inpainting.
-Combined together, MISATO contains 2000 examples from four inpainting domains, indoor, outdoor landscape, building, and background.
-As shown in Fig.
-6(b), we adopt the masking strategy as in Sec.
-3.1 but excluding the rectangle and complement rectangle masks.
+Combined together, MISATO contains 2000 examples from four inpainting domains, indoor, outdoor landscape, building, and background, as shown in Fig.
+10.
+we adopt the masking strategy as in Sec.
+3.1.1 but excluding the rectangle and complement rectangle masks.
 The masking ratio is set as $[0.2,0.8]$.
-5 Experiments
-Implementation details.
-ASUKA utilizes the stable diffusion text-guided inpainting model v1.5 as the generative backbone.
-We train ASUKA on Places2 [61].
-For the MAE [19] used in ASUKA, we train on images of size $256^{2}$, which is efficient and produce context-stable guidance for SD to generate high-resolution images.
-We fine-tune from the pre-trained MAE with a batch size of 1024.
-We use four self-attention blocks for the alignment module, training from scratch with AdamW [31] of learning rate 5e-2 with the standard diffusion objective.
+Figure 10: Different image domains in MISATO.
+Figure 11: The curse of self-attention, causing the MAE falsely estimate the masked region and powerful text-guided diffusion models fail to generation content based on text prompts.
+ASUKA potential circumvents this issue by using a blank paper image as the input to the MAE to provide correct prior.
+8 Implementation Details
+We use Places2 [101] to train ASUKA.
+For the MAE [31] used in ASUKA, we train on images of size $256^{2}$, which is efficient and produce context-stable guidance for generative models to generate high-resolution images.
+We fine-tune the MAE with a batch size of 1024.
+We train the alignment module with AdamW [53] of learning rate 5e-2 with the standard diffusion objective.
 We set $p$ as $100\%$ and linearly decay it to $10\%$ in the first 2K training steps and then freeze.
-We fine-tune the decoder based on [62] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay.
+For SD’s decoder, we fine-tune from [103] for 50K steps with a batch size of 40 and learning rate of 8e-5 with cosine decay.
+For FLUX’s decoder, we fine-tune from the original decoder with the same setup.
 We use ColorJitter for color augmentation, with brightness 0.15, contrast 0.2, saturation 0.1, and hue 0.03.
-Evaluation metrics.
-We use the Learned Perceptual Image Patch Similarity (LPIPS) [58] to calculate the patch-level image distances, Fréchet Inception Distance (FID) [20] to compare the distribution distance between generated images and real images, and Paired/Unpaired Inception Discriminative Score (P-IDS/U-IDS) [59] to measure the human-inspired linear separability.
-Competitors.
-We consider three variants of SD [39]: i) SD, prompted by "", performing unconditional generation; ii) SD-text, prompted by "background" as we have no captions in inpainting; iii) SD-prompt, prompted by learnable tokens which are trained using the same training pipeline of ASUKA.
-We also compare with state-of-the-art inpainting algorithms Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7], and SD-Repaint [32].
-Table 1: Quantitative comparison on MISATO and Places 2.
-Co-Mod [59], MAT [25], LaMa [45], MAE-FAR [7] and SD-Repaint [32] are state-of-the-art inpainting methods.
-SD [39] performs unconditional generation.
-SD-text uses “background” text prompt to guide generation.
-SD-prompt uses learnable prompts trained specifically for inpainting, using the same training setting as ASUKA, performing prompt-guided generation.
-ASUKA and SD variants use the stable diffusion text-guided inpainting model v1.5.
-Dataset MISATO (2k images) Places 2 (36.5k images)
+9 Further Analysis
+Table 5: Comparison of ASUKA with text-guided SD
 […]
-Co-Mod [59] 0.179 17.421 0.243 0.109 0.267 5.794 0.274 0.096
-MAT [25] 0.176 17.261 0.255 0.122 0.202 3.765 0.348 0.195
-LaMa [45] 0.155 15.436 0.260 0.135 0.202 6.693 0.247 0.050
-MAE-FAR [7] 0.142 13.283 0.282 0.153 0.174 3.559 0.307 0.105
-SD [39] 0.168 12.812 0.345 0.211 0.193 1.514 0.375 0.207
-SD-text 0.164 12.603 0.337 0.207 0.191 1.506 0.373 0.202
-SD-prompt 0.160 12.517 0.331 0.204 0.189 1.477 0.390 0.234
-SD-Repaint [32] 0.227 27.861 0.016 0.007 0.251 12.466 0.217 0.045
-ASUKA 0.150 11.460 0.368 0.256 0.183 1.230 0.413 0.287
-5.1 Comparison on Benchmarks
-Quantitative comparison results are reported in Tab.
-1.
-ASUKA, although based on frozen SD, achieves significant and consistent improvements over SD across all the evaluation metrics, and enjoys the state-of-the-art results on FID, U-IDS and P-IDS.
-Particularly, the U-IDS and P-IDS coincide with human preference [59], and has a potential upper bound of $0.5$, indicating the superiority of ASUKA by a large margin.
-The inconsistency results between LPIPS and FID, U-IDS, P-IDS, partially comes from the frozen SD, where ASUKA achieves consistent improvements but is bounded to the frozen U-Net.
-In summary, the results validate the contribution of ASUKA that yields context-stable and visual-consistent inpainting through frozen SD.
-This improvement is both validated in in-distribution dataset Places 2 and out-of-distribution dataset MISATO.
-Qualitative comparison examples are shown in Fig.
-8.
-i) The state-of-the-art inpainting algorithms usually suffer from unnatural generation, for example the unnatural boundaries in the third and fourth rows, and failed inpainting of tower in the third-to-last row.
-LaMa and MAE-FAR sometimes lead to blurred inpainting results, especially in the scenario of large continuous masks.
-ii) The SD variants usually suffer from the context-stability issue and hallucinate unreasonable objects, in almost all the illustrated images.
-iii) Particularly, for the background mask shown in the second-to-last row, state-of-the-art inpainting algorithms cannot identify the object precisely, and extend the human in hairs and elbows.
-SD hallucinates random elements in the masked region.
-iv) In contrast, ASUKA enjoys context-stable and visual-consistent inpainting.
-5.2 Further Analysis of ASUKA
-Ablation of MAE prior.
+SD (BLIP2) 0.163 12.536 0.370 0.225 0.880 70.846
+ASUKA-SD 0.150 11.495 0.423 0.312 0.958 47.753
+Comparison with text-guided inpainting
+We compare ASUKA with text-guided SD model, as shown in Tab.
+5.
+We run SD inpainting sing text captions generated by BLIP2 [43].
+ASUKA performs better, since captions describe the entire image, while MAE focuses on reconstructing only the masked region, leading to more precise guidance.
+Table 6: Ablation of $p$
+[…]
+p=0 0.155 11.804 0.403 0.288 0.940 48.032
+p=1 0.152 11.734 0.394 0.296 0.947 47.997
+linear decay p 0.152 11.558 0.405 0.307 0.955 47.814
+Ours 0.150 11.495 0.423 0.312 0.958 47.753
+Ablation of $p$
+We analyze how different values of $p$ affect ASUKA in Tab. 6. The results show that our warm-up and freeze strategy outperforms other approaches.
+Table 7: Additional results on benchmark datasets
+[…]
+CelebA-HQ SD 0.132 11.968 0.282 0.101 0.939 42.870
+ASUKA-SD 0.129 10.190 0.293 0.134 0.941 40.503
+FFHQ SD 0.139 2.235 0.371 0.197 0.944 43.529
+ASUKA-SD 0.131 2.060 0.386 0.205 0.955 30.848
+Additional Results
+We further compare ASUKA with standard SD on two additional datasets: CelebA-HQ [36] and FFHQ [37]. As shown in Tab. 7, these results provide more evidence of ASUKA’s effectiveness.
+Table 8: Our Decoder in Text-Guided Inpainting.
+[…]
+SD 0.297 0.180 30.255 0.312 0.930 57.136
+ASUKA-SD 0.298 0.175 29.350 0.350 0.931 38.123
+Our Decoder in Text-Guided Inpainting
+To test the generalizability of our decoder, we evaluate it on text-guided inpainting tasks.
+We compare our decoder with the original SD decoder using 1,000 randomly sampled images from “jackyhate/text-to-image-2M” [105].
+The results in Tab.
+8 confirm its effectiveness for general inpainting tasks.
+Table 9: Effect of each module.
+[…]
+SD w/ MAE 0.157 12.093 0.397 0.236 0.953 62.845
+SD w/ decoder 0.159 12.075 0.411 0.283 0.954 49.376
+ASUKA-SD 0.150 11.495 0.423 0.312 0.958 47.753
+Ablation on independent modules
+To understand the contribution of each module in ASUKA, we evaluate SD with the proposed modules added separately. The results, shown in Tab. 9, highlight the effectiveness of each module.
+Table 10: Comparison of ASUKA using pre-trained MAE v.s. fine-tuned MAE.
+[…]
+pre-trained 0.151 11.513 0.354 0.258
+fine-tuned 0.150 11.460 0.368 0.256
+Ablation of MAE prior
 We compare our fine-tuned MAE with directly adopting the MAE trained in [10].
 To this end, we train ASUKA with the MAE in [10] using the same training strategy and compare the results in Tab.
-2(a).
+10.
 Results suggest the improvements of fine-tuning MAE, especially on FID and U-IDS.
 This improvement comes from the better adaptation on the real-world masks.
-Ablation of alignment module.
-Our alignment module includes a dimension and a distribution alignment.
-We validate the effectiveness of each module by comparing with the following variants step by step: i) linear: Use linear layer to align feature dimension only; ii) attn: Based on i, further use a single self-attention block to align the distribution; iii) cross x4: we instead use learnable query and 4 cross-attention layers to learn the MAE prior.
-ASUKA adopts 4 self-attention blocks.
-Results are shown in Tab.
-2(b).
-The self-attention block shows improved results compared with only align dimension and cross-attention block.
-Using 4 self-attention blocks improves the capacity of the alignment module.
-Ablation of decoder.
-For the decoder, we compare ASUKA with i) VAE: the decoder used in SD; ii) + cond.: the decoder conditioned on unmasked image [62]; iii) + color: only use color augmentation to train the decoder; Results are in Tab.
-2(d), showing the superiority of our augmented decoder.
-Table 2: Ablation stuides of ASUKA.
-(a) Comparison of ASUKA using pre-trained (p.t.) MAE v.s. fine-tuned (f.t.) MAE.
-[…]
-p.t. 0.151 11.513 0.354 0.258
-f.t. 0.150 11.460 0.368 0.256
-(b) Ablation of different alignment modules. Linear adopts linear layer; attn adopts a self-attention layer; cross x4 adopts 4 cross-attention layers; ASUKA adopts 4 self-attention layers.
-[…]
-linear 0.155 11.934 0.361 0.227
-attn 0.152 11.613 0.362 0.234
-cross x4 0.152 11.762 0.368 0.238
-ASUKA 0.150 11.460 0.368 0.256
-(c) User-study of top-1 ratio among all the inpainting results.
-Context-stability (C.S.) measures the coherence between masked region and unmasked surroundings, with a preference of not generating new elements; Visual-consistency (V.C.) measures the color consistency.
-Model C.S.(%) V.C.(%)
+Table 11: User-study of top-1 ratio among all the inpainting results.
+Model UOM (%) CC(%)
 Co-Mod [99] 3.98 4.98
@@
 SD-Repaint [54] 1.60 0.55
-ASUKA 39.43 40.05
-(d) Comparison of different decoders for SD.
-VAE [39] is the original decoder used by SD; + cond.
-[62] is the decoder conditioned on unmasked image; + color uses the color augmentation strategy to perform local harmonization task; Ours further combines latent augmentation strategy to handle the gap between generated latent and real latent.
-[…]
-VAE 0.156 11.949 0.343 0.208
-+ cond. 0.151 11.634 0.361 0.231
-+ color 0.152 11.603 0.357 0.229
-Ours 0.150 11.460 0.368 0.256
-User-Study.
+ASUKA-SD 39.43 40.05
+User-study
 To evaluate the user preference on inpainting algorithms, we conduct an user-study.
 Specifically, we randomly select 40 testing images.
-We ask the user to select the best inpainting results from the following perspectives respectively: i) Context-stability (C.S.): the generated region should be context-stable with surrounding unmasked region, with a preference of not generating new elements; ii) Visual-consistency (V.C.) : the visual consistency between masked and unmasked regions.
+We ask the user to select the best inpainting results from the following perspectives respectively: i) Unwanted-object-mitigation (UOM): the generated region should be context-stable with surrounding unmasked region, with a preference of not generating new elements; ii) Color-consistency (CC) : the color consistency between masked and unmasked regions.
 We collect 100 valid anonymous questionnaire results, and report the average selection ratio among all the inpainting algorithms in Tab.
-2(c).
+11.
 This result validate the efficacy of ASUKA on alignment with human preference.
-Figure 9: The curse of self-attention, causing the MAE falsely estimate the masked region and powerful text-guided diffusion models fail to generation content based on text prompts.
-ASUKA potential circumvents this issue by using a blank paper image as the input to the MAE to provide correct prior.
-Limitation: The "curse" of self-attention.
+Limitation: The "curse" of self-attention
 The primary limitation of ASUKA stems from the inefficacy of the MAE prior, mainly due to issues within the self-attention module.
 Specifically, as shown in Fig.
-9, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal.
+11, the presence of multiple similar objects in an image may lead the MAE to incorrectly predict a similar object in the masked region, conflicting with the goal of object removal.
 Notably, this curse of self-attention significantly impacts diffusion-based generative models.
@@
 A more comprehensive solution would involve extra control on self-attention layers in diffusion models, which we leave as future work.
-6 Conclusion
-In this paper, we proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) to achieve context-stable and visual-consistent inpainting via frozen Stable Diffusion inpainting model (SD).
-To achieve context-stability, we adopt a reconstruction-based masked auto-encoder (MAE) as the context-stable prior for masked region purely from unmasked region.
-Then we align the context-stable prior to frozen SD with the proposed alignment module.
-To achieve visual-consistency, we resolve the mask-unmask color inconsistency of SD in the decoding process.
-We train an unmask-region conditioned KL-VAE decoder to perform local harmonization during the decoding process.
-To validate the efficacy of inpainting algorithms in different image domains and mask types, we introduce an evaluation dataset, named as MISATO, from existing datasets.
-ASUKA enjoys context-stable and visual-consistent inpainting results and superior than state-of-the-art inpainting algorithms and SD-based variants.
+Potential negative impact
+As an image editing tool, our proposed ASUKA will generate images based on user intentions for masking specific parts of the image, potentially resulting in unrealistic renderings and posing a risk of misuse.